### v1

# TSkips: Efficiency Through Explicit Temporal Delay Connections in Spiking Neural Networks

Abstract
Spiking Neural Networks (SNNs) with their bio-inspired Leaky Integrate-and-Fire (LIF) neurons inherently capture temporal information. This makes them well-suited for sequential tasks like processing event-based data from Dynamic Vision Sensors (DVS) and event-based speech tasks. Harnessing the temporal capabilities of SNNs requires mitigating vanishing spikes during training, capturing spatio-temporal patterns and enhancing precise spike timing. To address these challenges, we propose TSkips, augmenting SNN architectures with forward and backward skip connections that incorporate explicit temporal delays. These connections capture long-term spatio-temporal dependencies and facilitate better spike flow over long sequences. The introduction of TSkips creates a vast search space of possible configurations, encompassing skip positions and time delay values. To efficiently navigate this search space, this work leverages training-free Neural Architecture Search (NAS) to identify optimal network structures and corresponding delays. We demonstrate the effectiveness of our approach on four event-based datasets: DSEC-flow for optical flow estimation, DVS128 Gesture for hand gesture recognition and Spiking Heidelberg Digits (SHD) and Spiking Speech Commands (SSC) for speech recognition. Our method achieves significant improvements across these datasets: up to 18% reduction in Average Endpoint Error (AEE) on DSEC-flow, 8% increase in classification accuracy on DVS128 Gesture, and up to  8% and  16% higher classification accuracy on SHD and SSC, respectively.
1 Introduction
Figure 1: Model size vs. average endpoint error (AEE) for fully spiking EV-FlowNet architectures [29] modified with Forward and Backward TSkips on the DSEC-flow dataset (lower AEE is better). Each tick on the horizontal axis represents a 3×3$\times$3 × increase in model size. TSkips achieve an average reduction of 15% in AEE and 40.75% in inference energy (E) while maintaining comparable AEE to larger baselines.
Spiking Neural Networks (SNNs) are a class of bio-physically realistic models with Leaky Integrate-and-Fire (LIF) [1] neurons that offer a promising alternative for processing sequential data. By encoding information in the precise timing of spikes, SNNs capture spatio-temporal dependencies [43, 7] through the membrane potential of their LIF neurons. This intrinsic ability to process temporal information allows SNNs to excel at capturing dynamic patterns, effectively functioning as specialized RNNs [44] with reduced complexity.
The ability of SNNs to leverage the precise timing of spikes offers a distinct advantage over traditional Artificial Neural Networks (ANNs) for sequential tasks. Note, ANNs are “stateless” in nature [15] and their reliance on specially curated input encoding schemes [59, 38] hinder their ability to accurately capture precise timing information crucial in event streams, as generated by Dynamic Vision Sensors (DVS) [33, 4]. Standard Recurrent Neural Networks (vRNNs) [47] and Long Short-Term Memory (LSTM) networks [19, 40] attempt to address this, but remain difficult to train [39] and computationally expensive. Furthermore, transformers and spiking transformers [31, 57] excel at sequential processing but come at the cost of significantly larger models and computational overhead.
Deep SNNs offer the appealing combination of rich temporal information processing and computational efficiency, but training them effectively presents unique challenges [45]. The combination of multiple time steps with Back Propagation through Time (BPTT), non-differentiable spikes and sparse event data can worsen the vanishing spike problem [6]. To circumvent this, previous works have proposed adaptive LIF neurons with trainable threshold and leak [29] and approximate surrogate gradients [11, 35]. Other works [50, 52, 16] have considered synaptic delays as an additional parameter alongside weights, thus offering a richer representation of the spike patterns present in event data. We posit that these learnable synaptic delays only capture spatio-temporal patterns over a short window of time. Note, however, event data for speech recognition and optical flow also have long-term temporal dependencies that need to be captured through better spike representation.
In contrast, this work introduces explicit temporal delays in forward and backward skip connections, named temporal skips (TSkips), for SNNs and hybrid ANN-SNN [36] models. TSkips offer finer control over spike timing, enhance responsiveness to temporal patterns, mitigate vanishing spikes, all while capturing long-term spatio-temporal dependencies for event-based data. In addition, our method utilizes adaptive LIF neurons [29] that have learnable leaks and thresholds that capture local temporal patterns.
The next question to address would be, what are the optimal TSkip configurations within the network architecture and what are the corresponding temporal delays associated with them? To efficiently identify optimal TSkip configurations across different architectures and sequence lengths, we leverage training-free Neural Architecture Search (NAS) tailored for SNNs [27]. TSkips introduces minimal increase in model size with very few additional trainable parameters, minimal overhead and improves inference on event-based tasks. In addition, we have observed that training time is faster for the proposed networks compared to standard SNNs, RNNs and LSTMs. The contributions of this work can be summarized as follows:
  • We introduce TSkips, a novel mechanism that enhances SNNs and hybrid ANN-SNN architectures by enabling direct transmission of spike information between non-adjacent layers with temporal delays, effectively capturing long-term spatio-temporal patterns. (Section 3)
  • We analyzed various TSkip configurations, revealing the complexities of the search space, particularly the interplay between the temporal delay $(\Delta t)$, TSkip position and network depth. This complexity motivated the use of NAS to efficiently identify optimal TSkip architectures. (Section 4.2)
  • We show that TSkips achieves strong performance across sequential tasks. On the DSEC-flow dataset, they reduce average endpoint error (AEE) and inference energy, up to 18% and 56%, respectively, as shown in Fig. 1. Furthermore, TSkips achievess significant accuracy improvements on DVS128 Gesture, SHD and SSC datasets — up to 8.3%, 8.6% and 16.04%, respectively. This highlights the scalability and effectiveness of our approach across diverse sequential tasks and architectures. (Section 4)
2 Related Work
This section explores the challenges of training deep SNNs and incorporating delays to capture richer temporal dynamics.
2.1 Training Deep SNNs
Early methods to train SNNs relied on ANN-to-SNN conversion [6, 37]. However, this approach often struggled to match ANN performance on complex sequential tasks [10], particularly with event data. A key obstacle in training deep SNNs is the non-differentiable nature of spike activations [11], which prevents the direct application of gradient-based optimization techniques, leading to subpar performance.
To overcome this, approximate surrogate gradients [11, 35] have been proposed, that replace non-differentiable spike activations with smooth gradient approximations, enabling the use of BPTT [53] to train deep SNNs. However, this approach introduces other challenges like vanishing gradients, especially with long sequences and sparse event data [6], and reduce accuracy compared to ANNs (due to approximate gradients). To address these challenges, previous work has investigated two main avenues: adaptive LIF neurons with learnable thresholds and leaks [29], and traditional skip connections (without delays) in SNNs [2]. While adaptive LIF neurons allow for dynamic adjustment and better capture of spatio-temporal patterns, and skip connections can improve accuracy and efficiency, neither approach fully addresses the challenges of representing and capturing complex temporal patterns. To achieve this, we propose TSkips, incorporating explicit temporal delays into the network architecture with forward and backward skip connections.
2.2 Delays in SNNs
The importance of precise timing information in event-based data, such as that from dynamic vision sensors, has motivated research into effectively capturing temporal dynamics in SNNs [43, 7]. Some works have introduced delays alongside the weights in SNNs to offer a richer representation of the spike patterns. For instance, [50] introduced a fixed delay in SNNs to improve learning of temporal complexities, [52] extended this by making these delays learnable and introducing a layer-specific maximum delay. [16] further refined this by making the per-layer maximum delay learnable and utilizing Dilated Convolutions with Learnable Spacings (DCLS) [25] to capture spatio-temporal patterns. However, learning these delays, as in [52, 16], introduces a new set of parameters with the same dimension as the weights, essentially doubling the number of trainable parameters. Additionally, these methods only capture local temporal patterns over a short window of time, due to their reliance on the dynamics of LIF neurons.
In contrast to these methods that rely on learning delays, our proposed method introduces explicit temporal delays in the form of forward and backward skip connections (TSkips) within the SNN architecture. The use of TSkips with adaptive LIF neurons effectively captures both local and long-term spatio-temporal patterns with minimal additional computational overhead.
3 Methodology
This section details our methodology, starting with the input representation and the neuron model employed in our SNNs, followed by the proposed method, our NAS search space, and the training process.
3.1 Input Representation and Sensors
For efficient computation, SNNs require data in the form of discrete spikes, mirroring communication in biological neurons. DVS [33] cameras, output asynchronous spikes representing real-time changes in pixel intensity. This eliminates redundant data transmission and achieves higher temporal resolution than traditional frame-based cameras or rate coding methods [6]. These cameras trigger events when the log intensity change at a pixel crosses a threshold ($\theta$[12], represented as $||log(I_{t})-log(I_{t-1})||\geq\theta$. Events are encoded in the Address Event Representation (AER) format as tuples $(x,y,t,p)$ — representing pixel coordinates, timestamp, and polarity. Similarly, event-based audio datasets [8] mimic human auditory spiking activity, with spikes represented as $(x,t)$ — denoting spike unit and timestamp.
3.2 Neuron Model
The fundamental computational unit of SNNs is the spiking neuron. We employ Leaky Integrate-and-Fire (LIF) [1] neurons for their biological plausibility and computational efficiency. The LIF neuron integrates input spike information over time, accumulating it in its membrane potential. This potential gradually decays, allowing controlled forgetting of less relevant information. The dynamics of the LIF neuron can be described by
$U_{l}^{t}=\lambda U_{l}^{t-1}+W_{l}O_{l-1}^{t}-V_{l}^{th}O_{l}^{t-1}$ (1)
where $U_{l}^{t}$ represents the membrane potential of the neurons in layer l at time step t, $\lambda$ is the leak factor, controlling the decay rate of the membrane potential. $W_{l}$ is the weight matrix for neurons in layer $l$ and $V_{l}^{th}$ is the voltage threshold of layer $l$.
Figure 2: LIF Neuron
When the membrane potential surpasses the voltage threshold, the neuron emits a spike. The first term in Eq. (1) denotes the leakage in the membrane potential, the second term computes the weighted summation of output spikes from layer $(l-1)$ and the third term denotes the reduction in membrane potential when an output spike is generated at layer $l$. The membrane potential is then reset, either to be zero (hard reset) or by subtracting $V_{l}^{th}$ (soft reset). Equation (2) describes the spike generation.
$Z_{l}^{t}=\frac{U_{l}^{t}}{V_{l}^{th}}-1,\quad O_{l}^{t}=\begin{cases}1,&\text{if }Z_{l}^{t}>0\\ 0,&\text{otherwise}\end{cases}$ (2)
Figure 3: Illustration of TSkips in an SNN. (a) A 3-layer SNN with forward and backward TSkips, annotated with their origin and destination layers and the delay $(\Delta t)$. (b) Unrolled SNN with $T=4$ with a forward TSkip and $\Delta t=3$ compared with a standard skip connection (no delay). (c) Unrolled SNN with $T=4$ with a backward TSkip and $\Delta t=3$ compared with a standard backward connection ($\Delta t=1$[47]
We use a hard reset for optical flow estimation and a soft reset for gesture and speech classification. Fig. 2 shows the LIF neuron dynamics.
3.3 Proposed Method
Efficiently training deep SNNs is often hindered by vanishing gradients [6], which impede accurate spike propagation across time steps. To address this, we propose a novel method that augments skip connections in SNNs by introducing explicit temporal delays ($\Delta t$) within the connections. We refer to these augmented skip connections, which can be forward, backward, or a combination of both, as TSkips. The duration of these explicit delays is constrained by $t<\Delta t<T$, where $t$ is the current time step and $T$ is the sequence length of the data. This constraint on $\Delta t$ ensures that the network is not receiving information from the future.
To illustrate the information flow in a neural network with TSkips, let us consider how they modify the output of a layer. Using $h_{l}^{t}$ to denote the input to a layer $l$ at time step $t$, a forward/backward TSkip from layer $l\pm k$ at time step $t-\Delta t$ can be represented as:
$h_{l}^{t}=f_{l}(h_{l-1}^{t}\oplus W_{s}h_{l\pm k}^{(t-\Delta t)})$ (3)
where $f_{l}$ is the affine function of layer $l$, $\oplus$ represents either concatenation [22] or element-wise addition [17] and $W_{s}$ is a fixed shortcut path matrix [18] that randomly selects channels in $h_{l\pm k}^{(t-\Delta t)}$ to match the dimensions of $h_{l}^{t}$. Eq. 3 demonstrates that the output of layer $l$ at time step t is now influenced by the output of layer $l\pm k$ at time step $(t-\Delta t)$. Here, $k$ represents the difference in layer indices between connected layers. For example, if $l=4$ and $k=2$, a forward TSkip connects layers 2 and 4 with delay $\Delta t$.
Our method, incorporating TSkips, is illustrated in Fig. 3. Fig. 3(a) shows a simple 3-layer SNN architecture, illustrating both forward and backward TSkips between adjacent layers. Each connection is annotated with the origin and destination layers and $(\Delta t)$. Fig. 3(b) depicts the temporal unrolling of the network over $T=4$ time steps. It contrasts a standard skip connection $(\Delta t=0)$ and a forward TSkip with $\Delta t=3$. Fig. 3(c) similarly illustrates a backward TSkip with $\Delta t=3$ and contrasts it with a standard backward connection that has a delay of $\Delta t=1$, typically seen in vRNNs [47]. This comparison emphasizes that TSkips allow for more flexibility and can thus capture longer temporal dependencies compared to standard RNNs. This flexibility comes with minimal overhead, as TSkips are identity mappings that introduces very few additional parameters and no computational complexity to the network.
TSkips enhance the ability of SNNs and hybrid models to process sparse event streams by facilitating the propagation of temporally relevant information across non-adjacent layers. By capturing these long-term dependencies, TSkips can improve the network’s ability to learn complex temporal patterns. Furthermore, the explicit temporal delays in TSkips offer finer control over spike timing within the network. This finer control enhances the network’s responsiveness to temporal patterns by allowing it to selectively integrate information from different time steps. Additionally, by enabling the network to access relevant information from the past, TSkips can mitigate the issue of vanishing spikes. To further enhance temporal representation, we introduce a learnable scaling factor, $\alpha$, that controls the ratio of data at the current time-step ($t$) and data from the delayed time step ($t-\Delta t$) within the TSkip. This is incorporated into the skip as shown in Eq. 4:
$h_{l}^{t}=f_{l}(h_{l-1}^{t}\oplus W_{s}(\alpha h_{l\pm k}^{t}+(1-\alpha)h_{l\pm k}^{(t-\Delta t)}))$ (4)
3.4 Exploring the search space
Identifying optimal models with TSkips is challenging due to the vast number of potential architectures and associated TSkip configurations. To manage this complexity, we initialize our search with a backbone network [5]. This backbone provides a starting point for exploration, reducing the number of possible configurations and allowing the search to focus on optimizing the specific parameters of our proposed method. However, significant complexity remains due to the numerous choices for $\Delta t$, connection placement, and network depth.
To illustrate this complexity, consider a 3-layer network with the possibility of forward and backward connections between any two layers (Fig. 4). This simple example results in $12$ possible combinations of skip connections. With a sequence length of $T=6$, this grows to $120$ possible delay combinations and $2^{120}$ total possible TSkip configurations. This combinatorial problem, grows exponentially with the sequence lengths, as each additional time step introduces new TSkip configurations.
Figure 4: Illustration of the NAS Search Space. (a) Example of possible TSkips combinations in a 3-layer network with $\Delta t=6$. (b) Top-ranked forward and backward TSkips with $\Delta t=4$ and $\Delta t=3$, respectively, found by [27]. For visual clarity, the figure depicts the TSkips spanning a single delay window $(\Delta t)$.
This large search space necessitates an efficient exploration method to identify optimal architectures. To address this, we leverage NASWOT-SAHD [27], a training-free NAS method tailored for SNNs. NASWOT-SAHD scores networks at initialization using the Sparsity Aware Hamming Distance (SAHD), which quantifies the dissimilarity in spike patterns between different layers. By favoring networks with diverse spiking activity, SAHD acts as a proxy for identifying optimal sub networks of the backbone architecture augmented with TSkips. Specifically, [27] performs a random search for architectures with optimal $\Delta t$, TSkip positions (origin and destination) and network depth.
3.5 Training with Explicit Temporal Delays
Training SNNs and hybrid models with gradient descent requires addressing the non-differentiable nature of spike activations [11] in LIF neurons. We employ adaptive LIF neurons [29] with learnable leaks and thresholds, and use the ArcTangent [11] surrogate gradient to approximate the derivative of the spike activation function. To capture precise spike timing and spatio-temporal dependencies, we unroll the network in time. This, combined with adaptive LIF neurons and TSkips, allows for explicit representation of the network’s temporal evolution, capturing both long-term and local temporal patterns. We then use back propagation through time (BPTT) [53] for gradient propagation and Batch Normalization Through Time (BNTT) [26] for improved training stability.
4 Experiments
4.1 Experimental Setup
DSEC-Flow
The DSEC-Flow dataset [13] consists of 24 challenging driving sequences with varying lighting conditions, fast motion, and occlusions. To assess our models on unseen data, we created a custom test set by splitting the training set, following the approach in [41]. To evaluate the accuracy of the predicted optical flow, we use average end point error (AEE) [58], which measures the average Euclidean distance between the predicted and ground truth flow vectors.
$\vspace{-1ex}AEE=\frac{1}{n}\sum_{n}{\lVert{(u,v)}_{pred}-{(u,v)}_{gt}\rVert}_{2}$ (5)
For experiments on DSEC-flow [13], we use a fully spiking [29] and hybrid [36] multi-scale encoder-decoder network inspired by EV-FlowNet [58] as our backbone architecture. We augment this architecture by integrating TSkips into the existing skip connections between the encoder and decoder blocks, as depicted in Fig. 5. Due to the inherent structure of our backbone, the search space for optimal TSkips configurations is relatively small, focusing primarily on selecting which existing skip connection to modify and determining the appropriate temporal delay ($\Delta t$).
Figure 5: Fully spiking EV-FlowNet [58] architecture with forward and backward TSkips. Each TSkip replaces a pre-existing skip connection between the encoder and decoder. Although shown together for conciseness, forward and backward connections are evaluated independently. In practice, when used together, they connect different encoder-decoder layers. This architecture can be modified to be a hybrid ANN-SNN model, by replacing the LIF neurons with ReLU activations.
DVS128-Gesture
The IBM DVS128 Gesture dataset [21] contains 1342 instances of 11 hand gestures, captured with a DVS128 [33] camera. For experiments on DVS128 Gesture, we use a ResNet18 [17] backbone architecture [5] and explore various configurations by modifying its basic blocks. This search includes modifying the input and output channels of each layer, kernel sizes, stride, and network depth. To integrate TSkips in the backbone, we remove the original ResNet18 skip connections and search for optimal TSkip placement and corresponding $\Delta t$.
For optical flow estimation and gesture recognition, we introduce a learnable scaling factor, $\alpha$, within the TSkip to further enhance temporal processing (detailed in Section. 3.3).
SHD and SSC
The SHD [8] dataset comprises 10,000 recordings of spoken digits (zero to nine in English and German), while the larger SSC [8] dataset contains 100,000 recordings of spoken words from various speakers. These datasets present challenges for audio classification, requiring the network to capture and process subtle temporal patterns within spike trains that encode spoken sounds. We focus on the top-one classification task for all 35 classes in SSC and all 20 classes in SHD. For our experiments on speech recognition, we use a fully connected Multi Layer Perceptron(MLP) [47] as our backbone architecture. We explore various MLP configurations by adjusting the input and output features of each layer, the overall network depth, and the placement and $\Delta t$ of TSkips. In these experiments, we omit $\alpha$ used in the EV-FlowNet and DVS128 Gesture TSkip architectures.
To improve generalization, we augment the training sets for DVS128 Gesture, SHD and SSC with channel jitter and random noise [49]. We use sequence lengths (T) of 10, 30, and 99 for DSEC-flow, DVS128 Gesture, and SHD/SSC, respectively. All models are trained on an NVIDIA A40 GPU using the Adam optimizer [28]. For DSEC-flow, we use a multi-step learning rate scheduler (initial rate: $10^{-3}$, scaled by 0.7 every 10 epochs) and the supervised mean squared error (MSE) loss [29, 36] for 200 epochs. For all other datasets, we use a Cosine Annealing Learning Rate Scheduler [34] (initial rate: 0.001, minimum rate: $5\times 10^{-6}$, updated every 10 iterations) for 100 epochs.
Details on the constraints on the search parameters and TSkip configurations for DSEC-flow, DVS128 Gesture, SHD, and SSC can be found in the Supplementary material.
4.2 Ablation Study
Figure 6: Ablation study results on the SHD dataset demonstrating the impact of varying the (a) temporal delay $(\Delta t)$, (b) TSkip position and (c) network depth on classification accuracy.
As detailed in Section 3.4, the search space of possible TSkips configurations is exponentially large. To effectively navigate this complex search space, we performed an ablation study on an 8-layer MLP baseline found using [27] on the SHD dataset. We varied the temporal delay ($\Delta t$), TSkip position and network depth using concatenation-based [22] TSkips for our analysis (see Supplementary for detailed baseline network configuration and comparisons with addition-based TSkips).
Impact of Temporal Delay $\mathbf{(\Delta t)}$:
We investigated the impact of varying $\Delta t$ for both forward and backward TSkips (Fig. 6(a)). Forward connections connected the input data to the second layer, while backward connections connected the final output to the seventh layer, with a delay of $\Delta t$. Varying $\Delta t$ significantly influenced accuracy, with larger delays generally leading to better classification. This highlights the importance of optimizing $\Delta t$ relative to the sequence length ($T$), as both excessively small and large delays can hinder performance.
Impact of TSkip Position:
Next, we investigated the impact of TSkip position on classification accuracy (Fig. 6(b)). Fixing $\Delta t=16$, we varied the destination layer of both forward and backward connections. Forward connections originated from the input data, while backward connections originated from the final output. Our analysis revealed that connecting TSkips to deeper layers consistently improved accuracy. This suggests that propagating temporal information to deeper layers is crucial for enhancing model performance. Note, the optimal placement varies based on forward or backward TSkip.
Impact of Network Depth:
We analyzed the impact of TSkips on models of varying depth (Fig. 6(c)), with forward connections originating from the input and backward connections originating from the final output with $\Delta t=16$. Our analysis revealed that deeper networks benefited more from forward connections, while shallower architectures favored backward connections. This suggests the optimal configuration of TSkips and network depth are intertwined.
These findings underscore the complex interplay between $\Delta t$, TSkip position (origin and destination layers), and network depth, which we set as our search parameters. For the results presented in the following sections, we leveraged [27] to identify optimal baseline and concatenation-based [22] TSkip architectures, constraining the architecture search by the number of parameters.
4.3 DSEC-Flow Results
Incorporating TSkips into the fully spiking and hybrid EV-FlowNet architectures [29, 36], as detailed in Section 4.1, significantly lowers AEE on the DSEC-flow dataset (Table 1). Across both fully spiking and hybrid architectures, TSkips consistently yielded an average reduction in AEE of 14% and 9.5%, respectively. Notably, these improvements are achieved with models 3×3$\times$3 × smaller than baselines with comparable AEE. While SOTA methods like E-RAFT [14] and E-FlowFormer [32] achieve 6% and 10% lower AEE, respectively, these reductions come at the cost of increased complexity. E-RAFT, for example, while being 3×3$\times$3 × smaller than our best model, relies on complex mechanisms such as gated recurrent units and iterative flow refinement. Similalry, E-FlowFormer employs a computationally expensive transformer architecture and is trained on both a custom dataset and DSEC-flow. These complexities likely lead to higher inference energy compared to our lightweight TSkip architectures. Importantly, our approach achieves comparable qualitative performance to baselines 3×3$\times$3 × larger (Fig. 7), highlighting its efficiency in enabling accurate temporal processing without increasing model complexity. Hyper parameters and a detailed analysis of inference energy for TSkips on DSEC-flow are in the Supplementary material.
(a) GT
(b) Mini - Baseline
(c) Mini - F Tskip
(d) Mini - B Tskip
(e) Base - Baseline
Figure 7: Qualitative results on the DSEC-flow dataset. (a) Ground Truth (GT) mask, (b) Mini (3.4M) SNN baseline model mask, (c) Mini SNN (3.4M) with forward (F) Tskip, (d) Mini SNN (3.4M) with backward (B) TSkip, and (e) Base (13M) SNN baseline model mask. As observed, the Mini model that is 3.8$\times$ smaller than Base model performs just as well qualitatively when TSkips are incorporated.
Table 1: AEE on DSEC-flow (lower is better). Best performing SNN and hybrid baselines and forward(F)/backward(B) TSkip models are highlighted in bold.
Architecture #Params (M) AEE
SNN Hybrid
EV-FlowNet [14] 13.04 2.32 -
LSTM-FlowNet [41] N/A [^1] N/A - not reported #Params 1.28 -
E-RAFT [14] 5.3 0.79 -
E-FlowFormer [32] N/A 0.76 -
Our Models
Base - Baseline 13.04 1.35 0.96
Base + F TSkip 1.12 0.86
Base + B TSkip 1.13 0.84
Mini - Baseline 3.41 1.65 1.11
Mini + F TSkip 1.44 1.02
Mini + B TSkip 1.46 1.01
Micro - Baseline 0.93 1.80 1.22
Micro + F TSkip 1.57 1.12
Micro + B TSkip 1.56 1.09
Nano - Baseline 0.27 2.17 1.47
Nano + F TSkip 1.86 1.37
Nano + B TSkip 1.77 1.35
Pico - Baseline 0.092 2.57 2.02
Pico + F TSkip 2.19 1.78
Pico + B TSkip 2.28 1.81
Table 2: Classification accuracy on DVS-Gesture. Comparisons between forward(F)/backward(B) TSkips and SOTA models. Best models in bold.
Method Base Model #Params (M) Accuracy (%)
SLAYER [50] SNN (8 layer) N/A 93.6
KLIF [23] VGG-11 N/A 93.75
DECOLE [24] Custom N/A 95.2
SNN-Skip [2] ResNet18 N/A 95.43
SCRNN [54] Custom N/A 96.59
STS-ResNet [48] ResNet18 N/A 96.7
Mamba-Spike [42] Mambda 6.1 97.8
Our Models
Baseline ResNet18 Backbone 0.30 88.75
vRNN [^2] Backward skips are added to all baseline layers with $\Delta~{}t=1$. 0.30 89.12
F TSkip 0.42 94.82
B TSkip 0.38 95.97
F + B TSkip 0.55 97.52
4.4 DVS128-Gesture, SHD and SSC Results
Using TSkips in ResNet18 and MLP backbone architectures, as detailed in Section 4.1, consistently improved the classification accuracy of SNNs across the DVS Gesture, SHD, and SSC datasets (Tables 2 and 3).
On the DVS Gesture dataset, incorporating TSkips resulted in a significant 8.37% improvement in accuracy over the baseline model, a convolutional SNN found by [27]. Importantly, our approach achieved this with significantly smaller models compared to SOTA, including a spiking transformer [42], demonstrating the efficiency of our proposed method. Furthermore, our method outperformed several other approaches, including spiking RNNs [54], methods incorporating learnable delays [50], and methods using specialized LIF neurons or spatio-temporal feature extraction [23, 48]. This highlights the effectiveness of our approach in capturing and leveraging temporal dependencies for improved gesture recognition.
On the SHD dataset, TSkips achieved an average accuracy gain of 8.6% over the 4-layer baseline (Baseline - 1) and 8.31% over the 8-layer baseline (Baseline - 2). Our approach also outperformed various recurrent architectures (vRNNs, spiking RNNs, and LSTMs) by substantial margins. Similarly, on the SSC dataset, TSkips yielded an average accuracy gain of 14% over Baseline - 1 and 16.04% over the 8-layer Baseline - 2, again surpassing the performance of recurrent architectures.
Table 3: Classification accuracy on SHD and SSC datasets. Comparisons between forward(F), backward(B) and a combination (F+B) of both TSkips against SOTA models and two baseline models: a shallow 4-layer MLP network (Baseline - 1) and a deep 8-layer MLP (Baseline - 2). Best models with a single and two TSkips in bold.
Method #Params (M) Accuracy (%)
SHD SSC SHD SSC
SOTA Models
LSTM-LIF [56] 0.14 0.11 88.91 63.46
SRNN [55] N/A N/A 90.4 74.2
DL256-SNN-DLoss [51] 0.14 - 92.56 -
SpikGRU [9] - 0.28 - 77
radLIF [3] 3.9 3.9 94.62 77.40
Spikingformer [57]333[57] was evaluated on SHD and SSC. [57] for learning rate, with minimal updates to the architecture. 1.98 1.99 82.68 72.43
DCLS-Delays [52] 0.21 2.5 95.07 80.69
Our Baseline MLPs
Baseline - 1 0.16 0.12 84.32 64.19
Baseline - 2 1.04 1.04 86.42 67.54
Our MLP Models with TSkips
vRNN [^2] Backward skips are added to all baseline layers with $\Delta~{}t=1$. - 1 0.16 0.12 68.5 71.2
F TSkip - 1 0.24 0.42 92.32 76.5
B TSkip - 1 0.19 0.24 93.64 79.87
F + B TSkip - 1 0.20 0.54 93.01 78.64
vRNN [^2] Backward skips are added to all baseline layers with $\Delta~{}t=1$. - 2 1.04 1.04 71.35 72.24
F TSkip - 2 1.12 1.08 94.15 78.98
B TSkip - 2 1.16 1.14 93.86 79.65
F + B TSkip - 2 1.28 1.42 94.73 80.23
While TSkips achieves accuracy within 0.34% of SOTA model DCLS-Delays [16] on SHD and 0.46% on SSC, it offers several advantages: applicability across various network architectures, as shown in Section 4.3, efficiency due to very few new trainable parameters, and strong performance. Notably TSkips achieves comparable accuracy to [16] on the SSC dataset with a 44% smaller model.
These improvements can be attributed to the ability of TSkips to capture long-term temporal dependencies, facilitated by the unrolled network structure, detailed in Section 3. Our method improves both the accuracy and convergence rate of the models by mitigating vanishing spikes and enabling more efficient gradient propagation during training. Analysis in the Supplementary material demonstrates faster convergence and lower energy consumption compared to standard SNNs and recurrent architectures. Additionally, we show that incorporatingTSkips in ANNs and hybrid models significantly improves gesture and speech classification.
5 Conclusion
In conclusion, this work demonstrates the potential of incorporating temporal delays in forward and backward skip connections across SNNs and hybrid architectures. Our proposed method, optimized through training-free NAS, enhances accuracy, temporal representation, and efficiency while maintaining small model sizes, as shown by our results on the DSEC-flow, DVS128 Gesture, SHD, and SSC datasets. Furthermore, our ablation studies reveal the importance of understanding the interplay between network depth, temporal delays, and connection placement in optimizing SNN performance. Overall, this work demonstrates a promising direction for the development of efficient and accurate SNN architectures, achieving competitive performance with reduced complexity, thus offering a compelling alternative for real-world applications requiring fast and accurate processing of temporal information.
Acknowledgment
This work was supported by, the Center for the Co-Design of Cognitive Systems (COCOSYS), a DARPA-sponsored JUMP 2.0 center, the Semiconductor Research Corporation (SRC), and the National Science Foundation.
Supplementary Material
6 Inference Energy Estimation
To validate the observation that TSkips exhibit reduced energy consumption compared to baseline SNNs and standard recurrent methods (vRNNs [47] and LSTMs [19]) for various tasks (Sections 4.3 and 4.4), we estimate inference energy. These estimations are made using the method from [46, 30], which considers the energy required for sparse accumulate (AC) operations in SNNs ($E_{AC}=0.9pJ$) and dense multiply-and-accumulate (MAC) operations in ANNs ($E_{MAC}=4.6pJ$) based on a 45nm CMOS technology [20]. We calculated the total inference energy ($E_{total}$) as:
$E_{total}=\begin{cases}\#OPS_{SNN}\times E_{AC},&\text{for SNNs}\\ \#OPS_{ANN}\times E_{MAC},&\text{for ANNs}\end{cases}$ (6)
where #$\#OPS_{SNN}$ and #$\#OPS_{ANN}$ represent the layer-wise synaptic connections in SNNs and ANNs, respectively, calculated as:
#$\#OPS_{SNN}$ $=T\times\mathcal{N}\times\mathcal{C}\times\mathcal{M},$
#$\#OPS_{ANN}$ $=\mathcal{N}\times\mathcal{C}.$
Here, $T$ is the sequence length, $\mathcal{N}$ is the number of neurons in the layer, $\mathcal{C}$ is the number of synaptic connections per neuron, and $\mathcal{M}$ is the average spike rate over $T$.
6.1 Inference Energy Ablation Study
Figure 8: Ablation study on the SHD dataset demonstrating the impact of varying the (a) temporal delay $(\Delta t)$, (b) TSkip position and (c) network depth on inference energy.
Building upon our ablation study in Section 4.2, we further investigated the impact of different TSkip configurations on inference energy, estimated using Eq. 6. Specifically, we explored the impact of varying temporal delay ($\Delta t$), TSkip position, and network depth on an 8-layer multilayer perceptron (MLP) baseline model with concatenation-based TSkips, found using the Neural Architecture Search (NAS) framework [27]. This analysis is crucial because TSkips increases the average spike rate within the network, potentially leading to higher inference energy consumption.
Impact of Temporal Delay ($\Delta t$):
We investigated the impact of varying $\Delta t$ for both forward and backward TSkips on inference energy (Fig. 8(a)). Forward connections originated from the input data and terminated at the second layer, while backward connections originated from the final output and terminated at the seventh layer. Our analysis revealed that varying $\Delta t$ significantly influenced inference energy. Larger delays resulted in lower energy consumption, likely because fewer (but more temporally relevant) spikes were propagated through the connections.
Impact of TSkip Position:
Next, we investigated the impact of varying TSkips placement on inference energy (Fig. 8(b)). With a fixed $\Delta t=16$, we varied the destination layer of both forward (originating from the input data) and backward (originating from the final output) connections. While no clear trend emerged, terminating both forward and backward TSkips at deeper layers generally resulted in lower inference energy consumption.
Impact of Network Depth:
Finally, we analyzed the impact of TSkips on inference energy across models of varying depth (Fig. 8(c)). Forward connections originated from the input, and backward connections originated from the final output, with $\Delta t=16$. Our analysis revealed that increasing the number of layers led to a significant increase in inference energy.
This analysis of inference energy with varying TSkip parameters further emphasizes the complex interplay between $\Delta t$, TSkip position, and network depth. It highlights the importance of using methods like [27] to find optimal network architectures and TSkip configurations. While [27] does not explicitly optimize for minimal energy consumption, it prioritizes networks with diverse spiking activity by quantifying the dissimilarity between spiking patterns. As a result, the identified configurations often utilize TSkips with larger $\Delta t$ values, enabling our networks to efficiently capture long-term patterns while maintaining low energy consumption.
7 Additional Results - DSEC-flow
Table 4: Comparison of inference energy $(E_{total})$ using Eq. 6 for fully spiking DSEC-flow models with and without forward (F) / backward (B) TSkips across all scales.
Model #Params (×$\times 10^{6}$) $(\mathcal{N})$ Variant #OPSSNN (×$(\times 10^{9})$ Spike Rate $(\mathcal{M})$ $E_{total}$ (mJ)
Base 13.04 Baseline 25.9 45.83 23.3
F Tskip 32.7 57.8 29.4
B Tskip 30.7 54.32 27.6
Mini 3.41 Baseline 9.71 45.15 8.75
F Tskip 11.7 54.23 10.5
B Tskip 11.4 53.21 10.3
Micro 0.93 Baseline 5.81 61.9 5.25
F Tskip 6.14 65.4 5.52
B Tskip 5.86 62.4 5.27
Nano 0.27 Baseline 2.67 55.55 2.40
F Tskip 2.81 58.47 2.53
B Tskip 2.71 56.3 2.44
Pico 0.092 Baseline 2.11 74.12 1.90
F Tskip 2.29 80.6 2.07
B Tskip 2.25 78.9 2.02
7.1 Inference Energy Analysis
As demonstrated in Section 4.3, TSkips reduce average endpoint error (AEE) without increasing model complexity or energy consumption for optical flow estimation on the DSEC-flow dataset. To further analyze this, we evaluated the inference energy of our fully spiking EV-FlowNet architectures augmented with TSkips using Eq. 6.
Table 4 demonstrates the effectiveness of incorporating TSkips into fully spiking EV-FlowNet architectures [29]. By efficiently utilizing temporal information, our approach achieves comparable AEE to baseline models 3×3$\times$3 × larger, while significantly reducing inference energy. This improvement stems from TSkips ability to enhance spike propagation with minimal added parameters. While adding TSkips increases the average spike rate, the resulting models achieve comparable AEE to larger baselines (Table 1) with substantially lower energy consumption. This highlights the potential of our method for accurate, robust, and efficient optical flow prediction.
7.2 Convergence Analysis
Figure 9: Convergence analysis of fully spiking [29] and hybrid [36] EV-FlowNet architectures on the DSEC-flow dataset. The plots compare the convergence behavior of baseline models (Base and Nano) with their smaller forward (F) and backward (B) TSkip-augmented counterparts (Mini and Pico, respectively).
As discussed in Section 3, incorporating TSkips can lead to easier training and faster convergence for both fully spiking [29]and hybrid [36] EV-FlowNet architectures. To validate this, we analyzed the convergence behavior of our DSEC-flow models across different scales. Figure 9 shows that both our largest (Base vs. Mini with TSkips) and smallest (Nano vs. Pico with TSkips) models exhibit faster convergence when incorporating TSkips, for both fully spiking and hybrid variants. This faster convergence can be attributed to TSkips ability to mitigate vanishing spikes and enhance gradient flow throughout the network, enabling quicker optimization and improved learning dynamics. Additionally, by facilitating the propagation of temporally relevant information, TSkips can help the network learn important temporal dependencies more efficiently, leading to faster convergence. This consistent trend across different model sizes highlights the benefits of TSkips in facilitating fast, efficient and effective training.
7.3 Hyper parameters for DSEC-flow
Table 5 specifies the TSkip configurations used to replace existing skip connections in the fully spiking [29] and hybrid [36] EV-FlowNet architectures. Here, “TSkip Position” indicates the targeted connection, with ‘1’ representing the longest skip (between the first encoder and last decoder blocks) and ‘3’ the shortest (between the last encoder and first decoder blocks), see Fig. 5.
Table 5: TSkip parameters used for fully spiking and hybrid DSEC-flow models.
Architecture SNN Hybrid
$\Delta t$ TSkip Pos $\Delta t$ TSkip Pos
Base + F TSkip 3 1 3 1
Base + B TSkip 4 1 5 1
Mini + F TSkip 3 1 4 2
Mini + B TSkip 3 2 3 1
Micro + F TSkip 4 2 3 2
Micro + B TSkip 3 3 3 3
Nano + F TSkip 3 1 3 2
Nano + B TSkip 4 2 3 2
Pico + F TSkip 4 1 3 1
Pico + B TSkip 3 2 4 2
8 Additional Results - DVS128 Gesture, SHD and SSC
8.1 Inference Energy Analysis
As demonstrated in Section 4.4,TSkips improves classification accuracy for speech recognition compared to standard recurrent models (vRNNs and LSTMs). Additionally, we state that TSkips also reduce inference energy compared to these models. This energy efficiency is further confirmed by the estimations presented in Table 6 for both shallow (4-layer) and deep (8-layer) MLP [47] networks on the SHD [8] dataset. These estimations were calculated using Eq. 6, which considers the energy consumed by both sparse accumulate (AC) operations in SNNs and dense multiply-and-accumulate (MAC) operations in ANNs.
For comparison, the vRNNs in Table 6 are our baseline SNNs with backward skip connections added to every layer with $\Delta t=1$. The LSTMs are implemented as ANNs, with the $0.2\times 10^{6}$ and $1.1\times 10^{6}$ parameter models having $700$ input channels and hidden sizes of $75$ and $400$, respectively.
The observed reduction in energy consumption can be attributed to several key factors. Firstly, TSkips introduce a limited number of additional synaptic connections $(\mathcal{C})$ compared to the densely connected nature of RNNs and LSTMs. Secondly, they promote sparse spiking activity by selectively propagating information across time steps, leading to lower energy consumption than the dense computations inherent in recurrent layers. Finally, the explicit temporal delays in TSkips enable the network to focus on the most relevant information from previous time steps, thereby reducing unnecessary computations and further improving energy efficiency.
Table 6: Inference energy $(E_{total})$ best-performing models on the SHD dataset. Energy is calculated using Eq. 6.
Model #Params (×$\times 10^{6}$) $(\mathcal{N})$OPS (×$(\times 10^{12})$ Spike Rate (×$\times 10^{4}$) $(\mathcal{M})$ $E_{\text{total}}$ (mJ)
Baseline - 1 0.16 0.0083 3.16 7.52
vRNN - 1 0.16 0.213 38.08 192.41
LSTM - 1 0.2 0.016 - 74.63
F TSkip -1 0.24 0.038 5.62 34.79
B TSkip - 1 0.19 0.022 4.95 20.37
F + B TSkip - 1 0.20 0.032 7.42 29.06
Baseline - 2 1.04 0.815 7.63 734.26
vRNN - 2 1.04 10.854 58.23 9768.88
LSTM - 2 1.1 1.196 - 5505.28
F TSkip - 2 1.12 1.204 8.21 1083.90
B TSkip - 2 1.16 1.213 8.23 1091.87
F + B TSkip - 2 1.28 1.761 9.13 1585.70
8.2 Convergence Analysis
As discussed in Sections 3 and 4.4, incorporating TSkips can lead to easier training and faster convergence compared to recurrent models like vRNNs and LSTMs, as demonstrated in Fig. 10 for a 4-layer MLP network on the SHD dataset. This accelerated convergence can be attributed to several factors: improved gradient flow by mitigating the vanishing gradient problem, efficient information propagation by facilitating the access to temporally relevant information, and reduced computational complexity compared to the dense connectivity of recurrent networks. These factors, combined with the ability of TSkips to capture long-term temporal dependencies, contribute to the improved training efficiency observed in our experiments.
Figure 10: Test accuracy convergence on the SHD dataset, comparing the best-performing models with TSkips (from Table 3) to a vRNN and an LSTM with $0.2\times 10^{6}$ parameters.
8.3 TSkips in ANNs and Hybrid Models
As discussed in Section 4.4, incorporating TSkips into ANNs and hybrid ANN-SNN models can lead to improved classification accuracy. Table 7 presents these results, demonstrating the effectiveness of TSkips in enhancing both ANN and hybrid model performance, surpassing the performance of vRNNs and LSTMs, respectively.
Table 7: Classification accuracy of ANN and Hybrid ANN-SNN models on the SHD and SSC datasets. We compare the 4-layer MLP baseline (Baseline-1) with and without Forward (F) and Backward (B) TSkips.
Method #Params (×$(\times 10^{6})$ Accuracy (%)
ANN Hybrid
SHD
Baseline 0.16 49.95 74.11
vRNN 0.16 60.51 67.45
F TSkip 0.24 69.08 84.54
B TSkip 0.19 73.45 86.02
F + B TSkip 0.20 73.25 85.74
LSTM 0.20 80.54 -
SSC
Baseline 0.12 55.41 59.61
vRNN 0.12 68.29 71.43
F TSkip 0.42 71.19 73.48
B TSkip 0.24 70.23 72.87
F + B TSkip 0.54 71.09 74.96
LSTM 0.20 72.3 -
This improvement can be attributed to the ability of TSkips to facilitate more effective temporal processing. By incorporating explicit temporal delays, these connections allow the network to access and integrate information from past time steps, which can be crucial for understanding complex temporal patterns. This enhanced temporal awareness enables the network to make more accurate predictions, leading to improved classification accuracy. Furthermore, TSkips can mitigate the vanishing spikes problem often encountered in deep networks, by providing more direct pathways for gradient flow during training. This improved gradient propagation can lead to more stable and efficient training, further contributing to the enhanced performance observed in ANNs and hybrid models.
8.4 Comparison between TSkip Operators
To further analyze the impact of different TSkip connection types, we compared the performance of concatenation-based and addition-based TSkips on the 4-layer and 8-layer baselines (Table 3). Table 8 demonstrates the superior performance of concatenation-based TSkips over addition-based on the baseline models. Concatenation-based TSkips consistently outperformed their addition-based counterparts. This can be attributed to the increased representational capacity of concatenation, which expands the feature space by combining information from the skip connection and the destination layer. This richer representation allows the network to learn more complex temporal patterns, particularly crucial in deep SNNs where information is encoded in sparse spike trains and membrane potentials. Furthermore, concatenation preserves information from both sources, preventing potential information loss that can occur with addition.
Table 8: Comparison of classification accuracy on SHD and SSC datasets with concatenation-based (C) and addition-based (A) Forward (F) and Backward (B) TSkips on a 4-layer (Baseline -1) and 8-layer (Baseline - 2) MLP network.
Method #Params (×$(\times 10^{6})$ Accuracy (%)
SHD SSC SHD SSC
Baseline - 1
Baseline 0.16 0.12 84.32 64.19
F TSkip - C 0.24 0.42 92.32 76.5
B TSkip - C 0.19 0.24 93.64 79.87
F + B TSkip - C 0.20 0.54 93.01 78.64
F TSkip - A 0.19 0.37 87.65 72.13
B TSkip - A 0.11 0.21 87.54 73.15
F + B TSkip - A 0.11 0.40 88.13 75.12
Baseline - 2
Baseline 1.04 1.04 86.42 67.54
F TSkip - C 1.12 1.08 94.15 78.98
B TSkip - C 1.16 1.14 93.86 79.65
F + B TSkip - C 1.28 1.42 94.73 80.23
F TSkip - A 0.99 1.05 86.39 72.15
B TSkip - A 0.97 1.01 87.32 75.53
F + B TSkip - A 0.98 1.14 89.45 77.12
8.5 Hyper parameters for DVS128 Gesture, SHD and SSC
Table 9 presents the top-ranked network architectures and TSkip configurations identified by [27], with corresponding results shown in Tables  2 and 3.
The table specifies the following for each configuration:
  1. 1.
    Network architecture:
    1. (a)
      DVS128 Gesture: $2\times 64\times 64$ represents the input dimensions (polarity 2, $64\times 64$ spike resolution) for all networks. Convolutional layers are represented concisely, for example, 3c80s1 denotes a layer with a $3\times 3$ kernel, 80 output channels, and a stride of 1.
    2. (b)
      SHD and SSC: Specifies the number of input channels for each MLP layer, including the final fully connected layer. Channels represented as $n*2$ indicate that concatenation-based TSkips were used, doubling the number of channels at that layer. This representation is used to highlight that the results shown in Table 8, which uses addition-based TSkips, are based on the same underlying network architectures.
  2. 2.
    TSkip parameters: The temporal delay ($\Delta t$) and the origin and destination layers $(from\rightarrow to)$ for Forward (F), Backward (B), and combined (F+B) TSkips.
  3. 3.
    Other hyper parameters:
    1. (a)
      Leak and threshold initialization values for the adaptive LIF neurons: 0.6 and 15, respectively.
    2. (b)
      Dataset-specific dropout rates: 0.4 for SHD and 0.2 for SSC.
Table 9: Optimal Parameters for Training MLP Networks with Temporal Skip Connections on SHD and SSC Datasets.
Variant Network Delay ($\Delta t)$ TSkip Position (from $\rightarrow$ to)
DVS128 Gesture
Baseline 2$\times$64$\times$64-3c80s1-3c80s1- 5c86s1-5c64s1-1c64s1-1c32s11 - -
F TSkip 2$\times$64$\times$64-3c83s1-3c83s1-5c110s1- 3c94s1-1c94s1-1c32s11 5 1 $\rightarrow$ 2
B TSkip 2$\times$64$\times$64-3c64s1-5c64s1- 5c64s1-5c76s1-1c76s1-1c32s11 5 4 $\rightarrow$ 3
F + B TSkip 2$\times$64$\times$64-3c122s1-5c122s1- 3c122s1-5c64s1-1c64s1+1c32s11 F: 6 B: 8 F: 1 $\rightarrow$ 3 B: 5 $\rightarrow$ 4
SHD
Baseline - 1 700-124-288-144-20 - -
F TSkip - 1 700-124*2-432-115-20 24 1 $\rightarrow$ 2
B TSkip - 1 700*2-124-115-96-20 14 4 $\rightarrow$ 1
F + B TSkip - 1 700*2-124-144-72*2-20 F: 18 B: 14 F: 2 $\rightarrow$ 4 B: 2 $\rightarrow$ 1
Baseline - 2 700-512-288-224- 192-896-288-20 - -
F TSkip - 2 700-532-288-524- 192-448*2-288-20 17 2 $\rightarrow$ 6
B TSkip - 2 700-321-558*2-334 292-452-328-20) 12 3 $\rightarrow$ 5
F + B TSkip - 2 700-321-228*2-334- 492*2-462-448-20 F: 10 B: 13 F: 2 $\rightarrow$ 5 B: 7 $\rightarrow$ 3
SSC
Baseline - 1 700-124-148-130-35 - -
F TSkip - 1 700-324-298*2-145-35 15 1 $\rightarrow$ 3
B TSkip - 1 700-184*2-179-230-35 12 3 $\rightarrow$ 2
F + B TSkip - 1 700-364*2-257*2-195-35 F: 14 B: 8 F: 1 $\rightarrow$ 3 B: 4 $\rightarrow$ 2
Baseline - 2 700-731-410-340- 150-184-58-35 - -
F TSkip - 2 700-646-410-400- 250-184*2-135-35 17 2 $\rightarrow$ 6
B TSkip - 2 700-663-373*2-337- 292-182-135-35 15 7 $\rightarrow$ 3
F + B TSkip - 2 700-674-493*2-347*2- 278-192-89-35 F: 9 B: 11 F: 1 $\rightarrow$ 2 B: 7 $\rightarrow$ 3
9 NASWOT-SAHD Correlation on TSkips
To validate the effectiveness of NASWOT-SAHD [27] in identifying optimal network configurations, we analyzed the correlation between the accuracy and score of the identified networks, as shown in Fig. 11. The scores are based on the Sparsity Aware Hamming Distance (SAHD) [27], and the correlation is measured using Kendall’s $\tau$. Our baseline models achieved a $\tau$ correlation of 0.63, consistent with the results reported in [27]. However, the forward and backward TSkip network searches resulted in slightly lower correlations of 0.58 and 0.55, respectively. All results reported are for the top-ranked networks identified by  [27].
To ensure a focused and efficient search, we introduced several constraints:
  1. 1.
    Temporal Delay ($\Delta t$):
    1. (a)
      DSEC-flow ($T=10$): $2\leq\Delta t\leq 6$
    2. (b)
      DVS Gesture ($T=30$): $5\leq\Delta t\leq 14$
    3. (c)
      SHD and SSC ($T=99$): $10\leq\Delta t\leq 45$
  2. 2.
    TSkip Position: We constrained connections to not be within the same layer, ensuring that all connections originated from a different layer than the destination.
  3. 3.
    Model Size:
    1. (a)
      DSEC-Flow: No constraint was applied as we directly replaced existing skip connections in the fully spiking[29] and hybrid [36] EV-FlowNet architecture with TSkips, maintaining the original model size.
    2. (b)
      DVS128 Gesture: $0.6\times 10^{6}$ parameters
    3. (c)
      SHD and SSC (Baseline-1): $0.3\times 10^{6}$ parameters
    4. (d)
      SHD and SSC (Baseline-2): $1.3\times 10^{6}$ parameters
Figure 11: Correlation between accuracy and SAHD score for different network configurations (baseline, forward TSkip, and backward TSkip).

### v2

@@
 We demonstrate the effectiveness of our approach on four event-based datasets: DSEC-flow for optical flow estimation, DVS128 Gesture for hand gesture recognition and Spiking Heidelberg Digits (SHD) and Spiking Speech Commands (SSC) for speech recognition.
-Our method achieves significant improvements across these datasets: up to 18% reduction in Average Endpoint Error (AEE) on DSEC-flow, 8% increase in classification accuracy on DVS128 Gesture, and up to  8% and  16% higher classification accuracy on SHD and SSC, respectively.
+Our method achieves significant improvements across these datasets: up to 18% reduction in Average Endpoint Error (AEE) on DSEC-flow, 8% increase in classification accuracy on DVS128 Gesture, and up to $\sim 8\%$ and $\sim 16\%$ higher classification accuracy on SHD and SSC, respectively.
 1 Introduction
+Spiking Neural Networks (SNNs) are a class of bio-physically realistic models with Leaky Integrate-and-Fire (LIF) (Abbott, 1999) neurons that offer a promising alternative for processing sequential data.
+By encoding information in the precise timing of spikes, SNNs capture spatio-temporal dependencies (Rathi & Roy, 2020; Chowdhury et al., 2021; Wu et al., 2017) through the membrane potential of their LIF neurons.
+This intrinsic ability to process temporal information allows SNNs to excel at capturing dynamic patterns, effectively functioning as specialized RNNs (Rathi & Roy, 2024; Deng et al., 2022) with reduced complexity.
 Figure 1: Model size vs.
-average endpoint error (AEE) for fully spiking EV-FlowNet architectures [29] modified with Forward and Backward TSkips on the DSEC-flow dataset (lower AEE is better).
+average endpoint error (AEE) for fully spiking EV-FlowNet architectures (Kosta & Roy, 2023) modified with Forward and Backward TSkips on the DSEC-flow dataset (lower AEE is better).
 Each tick on the horizontal axis represents a 3×3$\times$3 × increase in model size.
 TSkips achieve an average reduction of 15% in AEE and 40.75% in inference energy (E) while maintaining comparable AEE to larger baselines.
-Spiking Neural Networks (SNNs) are a class of bio-physically realistic models with Leaky Integrate-and-Fire (LIF) [1] neurons that offer a promising alternative for processing sequential data.
-By encoding information in the precise timing of spikes, SNNs capture spatio-temporal dependencies [43, 7] through the membrane potential of their LIF neurons.
-This intrinsic ability to process temporal information allows SNNs to excel at capturing dynamic patterns, effectively functioning as specialized RNNs [44] with reduced complexity.
 The ability of SNNs to leverage the precise timing of spikes offers a distinct advantage over traditional Artificial Neural Networks (ANNs) for sequential tasks.
-Note, ANNs are “stateless” in nature [15] and their reliance on specially curated input encoding schemes [59, 38] hinder their ability to accurately capture precise timing information crucial in event streams, as generated by Dynamic Vision Sensors (DVS) [33, 4].
-Standard Recurrent Neural Networks (vRNNs) [47] and Long Short-Term Memory (LSTM) networks [19, 40] attempt to address this, but remain difficult to train [39] and computationally expensive.
-Furthermore, transformers and spiking transformers [31, 57] excel at sequential processing but come at the cost of significantly larger models and computational overhead.
-Deep SNNs offer the appealing combination of rich temporal information processing and computational efficiency, but training them effectively presents unique challenges [45].
-The combination of multiple time steps with Back Propagation through Time (BPTT), non-differentiable spikes and sparse event data can worsen the vanishing spike problem [6].
-To circumvent this, previous works have proposed adaptive LIF neurons with trainable threshold and leak [29] and approximate surrogate gradients [11, 35].
-Other works [50, 52, 16] have considered synaptic delays as an additional parameter alongside weights, thus offering a richer representation of the spike patterns present in event data.
-We posit that these learnable synaptic delays only capture spatio-temporal patterns over a short window of time.
-Note, however, event data for speech recognition and optical flow also have long-term temporal dependencies that need to be captured through better spike representation.
-In contrast, this work introduces explicit temporal delays in forward and backward skip connections, named temporal skips (TSkips), for SNNs and hybrid ANN-SNN [36] models.
+Note, ANNs are “stateless” in nature (Hagenaars et al., 2021b) and their reliance on specially curated input encoding schemes (Zhu et al., 2018b; Paredes-Vallés & De Croon, 2021) hinder their ability to accurately capture precise timing information crucial in event streams, as generated by Dynamic Vision Sensors (DVS) (Lichtsteiner et al., 2008; Brandli et al., 2014).
+Standard Recurrent Neural Networks (vRNNs) (Rumelhart et al., 1986) and Long Short-Term Memory (LSTM) networks (Hochreiter, 1997; Ponghiran & Roy, 2021; Gehrig et al., 2021b) attempt to address this, but remain difficult to train (Pascanu et al., 2013) and computationally expensive.
+Furthermore, transformers and spiking transformers (Li et al., 2022a; Zhou et al., 2023; Li et al., 2023) excel at sequential processing but come at the cost of significantly larger models and computational overhead.
+Deep SNNs offer the appealing combination of rich temporal information processing and computational efficiency, but training them effectively presents unique challenges (Pfeiffer & Pfeil, 2018; Rathi et al., 2021).
+The combination of multiple time steps with Back Propagation through Time (BPTT), non-differentiable spikes and sparse event data can worsen the vanishing spike problem (Hagenaars et al., 2021a; Neftci et al., 2019).
+To circumvent this, previous works have proposed adaptive LIF neurons with trainable threshold and leak (Fang et al., 2021b; Kosta & Roy, 2023; Wang & Li, 2023) and approximate surrogate gradients (Fang et al., 2021a; Neftci et al., 2019).
+Other works (Shrestha & Orchard, 2018; Sun et al., 2023b; Hammouamri et al., 2023) have considered synaptic delays as an additional parameter alongside weights, thus offering a richer representation of the spike patterns present in event data.
+In contrast, this work introduces explicit temporal delays in forward and backward skip connections, named temporal skips (TSkips), for SNNs and hybrid ANN-SNN (Kugele et al., 2021; Negi et al., 2024) models.
 TSkips offer finer control over spike timing, enhance responsiveness to temporal patterns, mitigate vanishing spikes, all while capturing long-term spatio-temporal dependencies for event-based data.
-In addition, our method utilizes adaptive LIF neurons [29] that have learnable leaks and thresholds that capture local temporal patterns.
-The next question to address would be, what are the optimal TSkip configurations within the network architecture and what are the corresponding temporal delays associated with them?
-To efficiently identify optimal TSkip configurations across different architectures and sequence lengths, we leverage training-free Neural Architecture Search (NAS) tailored for SNNs [27].
+In addition, our method utilizes adaptive LIF neurons (Yin et al., 2020; Hagenaars et al., 2021a; Kosta & Roy, 2023) that have learnable leaks and thresholds that capture local temporal patterns.
+The next question to address would be, what are the optimal TSkips configurations within the network architecture and what are the corresponding temporal delays associated with them?
+To efficiently identify optimal TSkips configurations across different architectures and sequence lengths, we leverage training-free Neural Architecture Search (NAS) tailored for SNNs (Kim et al., 2022).
 TSkips introduces minimal increase in model size with very few additional trainable parameters, minimal overhead and improves inference on event-based tasks.
@@
 (Section 3)
-  • We analyzed various TSkip configurations, revealing the complexities of the search space, particularly the interplay between the temporal delay $(\Delta t)$, TSkip position and network depth.
-This complexity motivated the use of NAS to efficiently identify optimal TSkip architectures.
+  • We analyzed various TSkips configurations, revealing the complexities of the search space, particularly the interplay between the temporal delay $(\Delta t)$, TSkips position and network depth.
+This complexity motivated the use of NAS to efficiently identify optimal TSkips architectures.
 (Section 4.2)
@@
 2.1 Training Deep SNNs
-Early methods to train SNNs relied on ANN-to-SNN conversion [6, 37].
-However, this approach often struggled to match ANN performance on complex sequential tasks [10], particularly with event data.
-A key obstacle in training deep SNNs is the non-differentiable nature of spike activations [11], which prevents the direct application of gradient-based optimization techniques, leading to subpar performance.
-To overcome this, approximate surrogate gradients [11, 35] have been proposed, that replace non-differentiable spike activations with smooth gradient approximations, enabling the use of BPTT [53] to train deep SNNs.
-However, this approach introduces other challenges like vanishing gradients, especially with long sequences and sparse event data [6], and reduce accuracy compared to ANNs (due to approximate gradients).
-To address these challenges, previous work has investigated two main avenues: adaptive LIF neurons with learnable thresholds and leaks [29], and traditional skip connections (without delays) in SNNs [2].
-While adaptive LIF neurons allow for dynamic adjustment and better capture of spatio-temporal patterns, and skip connections can improve accuracy and efficiency, neither approach fully addresses the challenges of representing and capturing complex temporal patterns.
+Early methods to train SNNs relied on ANN-to-SNN conversion (Cao et al., 2015; Panda et al., 2020).
+However, this approach often struggled to match ANN performance on complex sequential tasks (Deng et al., 2020), particularly with event data.
+To overcome the challenges associated with ANN-to-SNN conversion, Kugele et al.
+(2020) focus on converting existing neural network architectures into spiking neural networks using streaming rollouts (Fischer et al., 2018).
+By temporally unfolding the recurrent structures in the spiking domain the DenseNet skip connections essentially have a single step delay, much like vRNNs.
+A key obstacle in training deep SNNs is the non-differentiable nature of spike activations (Fang et al., 2021a), which prevents the direct application of gradient-based optimization techniques, leading to subpar performance.
+To overcome this, approximate surrogate gradients (Fang et al., 2021a; Neftci et al., 2019; Wu et al., 2017) have been proposed, that replace non-differentiable spike activations with smooth gradient approximations, enabling the use of BPTT (Werbos, 1990) to train deep SNNs.
+However, this approach introduces other challenges like vanishing gradients, especially with long sequences and sparse event data (Hagenaars et al., 2021a; Neftci et al., 2019), and reduce accuracy compared to ANNs (due to approximate gradients).
+To address these challenges, previous work has investigated two main avenues: adaptive LIF neurons with learnable thresholds and leaks (Fang et al., 2021b; Kosta & Roy, 2023; Wang & Li, 2023), and skip connections in SNNs (Benmeziane et al., 2023).
+While adaptive LIF neurons allow for dynamic adjustment and better capture of spatio-temporal patterns, skip connections can improve accuracy and efficiency, However, neither approach fully addresses the challenges of representing and capturing complex temporal patterns.
 To achieve this, we propose TSkips, incorporating explicit temporal delays into the network architecture with forward and backward skip connections.
 2.2 Delays in SNNs
-The importance of precise timing information in event-based data, such as that from dynamic vision sensors, has motivated research into effectively capturing temporal dynamics in SNNs [43, 7].
-Some works have introduced delays alongside the weights in SNNs to offer a richer representation of the spike patterns.
-For instance, [50] introduced a fixed delay in SNNs to improve learning of temporal complexities, [52] extended this by making these delays learnable and introducing a layer-specific maximum delay.
-[16] further refined this by making the per-layer maximum delay learnable and utilizing Dilated Convolutions with Learnable Spacings (DCLS) [25] to capture spatio-temporal patterns.
-However, learning these delays, as in [52, 16], introduces a new set of parameters with the same dimension as the weights, essentially doubling the number of trainable parameters.
-Additionally, these methods only capture local temporal patterns over a short window of time, due to their reliance on the dynamics of LIF neurons.
-In contrast to these methods that rely on learning delays, our proposed method introduces explicit temporal delays in the form of forward and backward skip connections (TSkips) within the SNN architecture.
-The use of TSkips with adaptive LIF neurons effectively captures both local and long-term spatio-temporal patterns with minimal additional computational overhead.
+The importance of precise timing information in event-based data, such as that from dynamic vision sensors, has motivated research into effectively capturing temporal dynamics in SNNs (Rathi & Roy, 2020; Chowdhury et al., 2021).
+SCTT (Soo et al., 2023) explore temporal skips in recurrent networks and are evalauted on cognitive tasks that have long-term temporal dependencies.
+However, SCTT employs an expensive evaluation strategy to identify optimal delay and temporal skip hyperparameters that involves training several models across multiple tasks to identify optimal connections.
+TTFS (Kim et al., 2024) explore adding delays in skip connections by encoding the latency of the first spike emitted by a neuron.
+This encoded delay is then introduced into skip connections in SNNs as learnable parameters that promote local path synchronization.
+However, the evaluation of both these works are limited to simple tasks that are not typically conducive to SNNs.
+Furthermore, while the spike encoding strategy of TTFS allows for better synchronization of spikes for the model, it does not account for factors such as noise or irregularities in the event stream or long periods of sparse events.
+Other works have introduced delays alongside the weights in SNNs to offer a richer representation of the spike patterns.
+For instance, Shrestha & Orchard (2018) introduced a fixed delay in SNNs to improve learning of temporal complexities, Sun et al.
+(2023b) extended this by making these delays learnable and introducing a layer-specific maximum delay.
+Hammouamri et al.
+(2023) further refined this by making the per-layer maximum delay learnable and utilizing Dilated Convolutions with Learnable Spacings (DCLS) (Khalfaoui-Hassani et al., 2023) to capture spatio-temporal patterns.
+However, learning these delays, as in Sun et al.
+(2023b); Hammouamri et al.
+(2023), introduces a new set of parameters with the same dimension as the weights, essentially doubling the number of trainable parameters.
+In contrast to these methods that rely on expensive tempral skip configuration evaluation strategies, estimating delay values for local path synchronization or learning delays, our proposed method introduces explicit temporal delays in SNNs.
+We explore these delays in both forward and backward temporal skip connections, named TSkips.
+The use of TSkips coupled with a learnable scaling factor and adaptive LIF neurons effectively captures long-term spatio-temporal patterns with minimal additional computational overhead.
 3 Methodology
@@
 For efficient computation, SNNs require data in the form of discrete spikes, mirroring communication in biological neurons.
-DVS [33] cameras, output asynchronous spikes representing real-time changes in pixel intensity.
-This eliminates redundant data transmission and achieves higher temporal resolution than traditional frame-based cameras or rate coding methods [6].
-These cameras trigger events when the log intensity change at a pixel crosses a threshold ($\theta$[12], represented as $||log(I_{t})-log(I_{t-1})||\geq\theta$.
+DVS (Lichtsteiner et al., 2008) cameras output asynchronous spikes representing real-time changes in pixel intensity $(I_{t})$.
+This eliminates redundant data transmission and achieves higher temporal resolution than traditional frame-based cameras or rate coding methods (Cao et al., 2015).
+These cameras trigger events when the log intensity change at a pixel crosses a threshold ($\theta$(Gallego et al., 2022), represented as $||log(I_{t})-log(I_{t-1})||\geq\theta$.
 Events are encoded in the Address Event Representation (AER) format as tuples $(x,y,t,p)$ — representing pixel coordinates, timestamp, and polarity.
-Similarly, event-based audio datasets [8] mimic human auditory spiking activity, with spikes represented as $(x,t)$ — denoting spike unit and timestamp.
+Similarly, event-based audio datasets (Cramer et al., 2022) mimic human auditory spiking activity, with spikes represented as $(x,t)$ — denoting spike unit and timestamp.
 3.2 Neuron Model
 The fundamental computational unit of SNNs is the spiking neuron.
-We employ Leaky Integrate-and-Fire (LIF) [1] neurons for their biological plausibility and computational efficiency.
+We employ Leaky Integrate-and-Fire (LIF) (Abbott, 1999) neurons for their biological plausibility and computational efficiency.
 The LIF neuron integrates input spike information over time, accumulating it in its membrane potential.
@@
 where $U_{l}^{t}$ represents the membrane potential of the neurons in layer l at time step t, $\lambda$ is the leak factor, controlling the decay rate of the membrane potential.
-$W_{l}$ is the weight matrix for neurons in layer $l$ and $V_{l}^{th}$ is the voltage threshold of layer $l$.
+$W_{l}$ is the weight matrix for neurons in layer $l$, $V_{l}^{th}$ is the voltage threshold of layer $l$ and $O_{l}^{t}$ is the output generated by layer $l$ at time step t.
 Figure 2: LIF Neuron
@@
 $Z_{l}^{t}=\frac{U_{l}^{t}}{V_{l}^{th}}-1,\quad O_{l}^{t}=\begin{cases}1,&\text{if }Z_{l}^{t}>0\\ 0,&\text{otherwise}\end{cases}$ (2)
+We use a hard reset for optical flow estimation and a soft reset for gesture and speech classification. Fig. 2 shows the LIF neuron dynamics.
+3.3 Proposed Method
 Figure 3: Illustration of TSkips in an SNN.
 (a) A 3-layer SNN with forward and backward TSkips, annotated with their origin and destination layers and the delay $(\Delta t)$.
-(b) Unrolled SNN with $T=4$ with a forward TSkip and $\Delta t=3$ compared with a standard skip connection (no delay).
-(c) Unrolled SNN with $T=4$ with a backward TSkip and $\Delta t=3$ compared with a standard backward connection ($\Delta t=1$[47]
-We use a hard reset for optical flow estimation and a soft reset for gesture and speech classification. Fig. 2 shows the LIF neuron dynamics.
-3.3 Proposed Method
-Efficiently training deep SNNs is often hindered by vanishing gradients [6], which impede accurate spike propagation across time steps.
+(b) Unrolled SNN with $T=4$ with a forward TSkips and $\Delta t=3$ compared with a standard skip connection (no delay).
+(c) Unrolled SNN with $T=4$ with a backward TSkips and $\Delta t=3$ compared with a standard backward connection ($\Delta t=1$(Rumelhart et al., 1986)
+Efficiently training deep SNNs is often hindered by vanishing gradients (Hagenaars et al., 2021a; Neftci et al., 2019), which impede accurate spike propagation across time steps.
 To address this, we propose a novel method that augments skip connections in SNNs by introducing explicit temporal delays ($\Delta t$) within the connections.
 We refer to these augmented skip connections, which can be forward, backward, or a combination of both, as TSkips.
-The duration of these explicit delays is constrained by $t<\Delta t<T$, where $t$ is the current time step and $T$ is the sequence length of the data.
-This constraint on $\Delta t$ ensures that the network is not receiving information from the future.
+The duration of these explicit delays is constrained by $0<t-\Delta t<T$, where $t$ is the current time step and $T$ is the sequence length of the data.
+Specifically, $t$ represents the set of all discrete time steps within the sequence, denoted as $t=\left\{t_{i},\forall i=0,\cdots,T\right\}$.
+This constraint on $\Delta t$ ensures that the network does not receive future information.
 To illustrate the information flow in a neural network with TSkips, let us consider how they modify the output of a layer.
-Using $h_{l}^{t}$ to denote the input to a layer $l$ at time step $t$, a forward/backward TSkip from layer $l\pm k$ at time step $t-\Delta t$ can be represented as:
+If $h_{l}^{t}$ denotes the input to a layer $l$ at time step $t$, then a forward TSkips provides input $h_{l-k}^{t}$ and a backward TSkips provides input $h_{l+k}^{t}$.
+For brevity, we denote these skip inputs as $h_{l\pm k}^{t}$, where $k$ indicates the layer separation between layer $l$ and the source of the skip connection.
+TSkips can be represented as:
 $h_{l}^{t}=f_{l}(h_{l-1}^{t}\oplus W_{s}h_{l\pm k}^{(t-\Delta t)})$ (3)
-where $f_{l}$ is the affine function of layer $l$, $\oplus$ represents either concatenation [22] or element-wise addition [17] and $W_{s}$ is a fixed shortcut path matrix [18] that randomly selects channels in $h_{l\pm k}^{(t-\Delta t)}$ to match the dimensions of $h_{l}^{t}$.
+where $f_{l}$ is the affine function of layer $l$, $\oplus$ represents either concatenation (Huang et al., 2017) or element-wise addition (He et al., 2015) and $W_{s}$ is a fixed shortcut path matrix (He et al., 2016) that randomly selects channels in $h_{l\pm k}^{(t-\Delta t)}$ to match the dimensions of $h_{l}^{t}$.
 Eq.
@@
 Here, $k$ represents the difference in layer indices between connected layers.
-For example, if $l=4$ and $k=2$, a forward TSkip connects layers 2 and 4 with delay $\Delta t$.
+For example, if $l=4$ and $k=2$, a forward TSkips connects layers 2 and 4 with delay $\Delta t$.
 Our method, incorporating TSkips, is illustrated in Fig.
@@
 3(b) depicts the temporal unrolling of the network over $T=4$ time steps.
-It contrasts a standard skip connection $(\Delta t=0)$ and a forward TSkip with $\Delta t=3$.
+It contrasts a standard skip connection $(\Delta t=0)$ and a forward TSkips with $\Delta t=3$.
 Fig.
-3(c) similarly illustrates a backward TSkip with $\Delta t=3$ and contrasts it with a standard backward connection that has a delay of $\Delta t=1$, typically seen in vRNNs [47].
+3(c) similarly illustrates a backward TSkips with $\Delta t=3$ and contrasts it with a standard backward connection that has a delay of $\Delta t=1$, typically seen in vRNNs (Rumelhart et al., 1986).
 This comparison emphasizes that TSkips allow for more flexibility and can thus capture longer temporal dependencies compared to standard RNNs.
@@
 Additionally, by enabling the network to access relevant information from the past, TSkips can mitigate the issue of vanishing spikes.
-To further enhance temporal representation, we introduce a learnable scaling factor, $\alpha$, that controls the ratio of data at the current time-step ($t$) and data from the delayed time step ($t-\Delta t$) within the TSkip.
+To further enhance temporal representation, we introduce a learnable scaling factor, $\alpha$, that controls the ratio of data at the current time-step ($t$) and data from the delayed time step ($t-\Delta t$) within the TSkips.
 This is incorporated into the skip as shown in Eq.
@@
 3.4 Exploring the search space
-Identifying optimal models with TSkips is challenging due to the vast number of potential architectures and associated TSkip configurations.
-To manage this complexity, we initialize our search with a backbone network [5].
+Figure 4: Illustration of the NAS Search Space.
+(a) Example of possible TSkips combinations in a 3-layer network with $T=6$.
+(b) Top-ranked forward and backward TSkips with $\Delta t=4$ and $\Delta t=3$, respectively, found by (Kim et al., 2022).
+For visual clarity, the figure depicts the TSkips spanning a single delay window $(\Delta t)$.
+Here $t_{0}$ refers to the first time step and $T$ is the sequence length.
+Identifying optimal models with TSkips is challenging due to the vast number of potential architectures and associated TSkips configurations.
+To manage this complexity, we initialize our search with a backbone network (Cai et al., 2020).
 This backbone provides a starting point for exploration, reducing the number of possible configurations and allowing the search to focus on optimizing the specific parameters of our proposed method.
@@
 This simple example results in $12$ possible combinations of skip connections.
-With a sequence length of $T=6$, this grows to $120$ possible delay combinations and $2^{120}$ total possible TSkip configurations.
-This combinatorial problem, grows exponentially with the sequence lengths, as each additional time step introduces new TSkip configurations.
-Figure 4: Illustration of the NAS Search Space.
-(a) Example of possible TSkips combinations in a 3-layer network with $\Delta t=6$.
-(b) Top-ranked forward and backward TSkips with $\Delta t=4$ and $\Delta t=3$, respectively, found by [27].
-For visual clarity, the figure depicts the TSkips spanning a single delay window $(\Delta t)$.
+With a sequence length of $T=6$, this grows to $120$ possible delay combinations and $2^{120}$ total possible TSkips configurations.
+This combinatorial problem, grows exponentially with the sequence lengths, as each additional time step introduces new TSkips configurations.
 This large search space necessitates an efficient exploration method to identify optimal architectures.
-To address this, we leverage NASWOT-SAHD [27], a training-free NAS method tailored for SNNs.
+To address this, we leverage NASWOT-SAHD (Kim et al., 2022), a training-free NAS method tailored for SNNs.
 NASWOT-SAHD scores networks at initialization using the Sparsity Aware Hamming Distance (SAHD), which quantifies the dissimilarity in spike patterns between different layers.
 By favoring networks with diverse spiking activity, SAHD acts as a proxy for identifying optimal sub networks of the backbone architecture augmented with TSkips.
-Specifically, [27] performs a random search for architectures with optimal $\Delta t$, TSkip positions (origin and destination) and network depth.
+Specifically, (Kim et al., 2022) performs a random search for architectures with optimal $\Delta t$, TSkips positions (origin and destination) and network depth.
 3.5 Training with Explicit Temporal Delays
-Training SNNs and hybrid models with gradient descent requires addressing the non-differentiable nature of spike activations [11] in LIF neurons.
-We employ adaptive LIF neurons [29] with learnable leaks and thresholds, and use the ArcTangent [11] surrogate gradient to approximate the derivative of the spike activation function.
+Training SNNs and hybrid models with gradient descent requires addressing the non-differentiable nature of spike activations (Fang et al., 2021a) in LIF neurons.
+We employ adaptive LIF neurons (Kosta & Roy, 2023) with learnable leaks and thresholds, and use the ArcTangent (Fang et al., 2021a) surrogate gradient to approximate the derivative of the spike activation function.
 To capture precise spike timing and spatio-temporal dependencies, we unroll the network in time.
 This, combined with adaptive LIF neurons and TSkips, allows for explicit representation of the network’s temporal evolution, capturing both long-term and local temporal patterns.
-We then use back propagation through time (BPTT) [53] for gradient propagation and Batch Normalization Through Time (BNTT) [26] for improved training stability.
+We then use back propagation through time (BPTT) (Werbos, 1990) for gradient propagation and Batch Normalization Through Time (BNTT) (Kim & Panda, 2021) for improved training stability.
 4 Experiments
@@
 DSEC-Flow
-The DSEC-Flow dataset [13] consists of 24 challenging driving sequences with varying lighting conditions, fast motion, and occlusions.
-To assess our models on unseen data, we created a custom test set by splitting the training set, following the approach in [41].
-To evaluate the accuracy of the predicted optical flow, we use average end point error (AEE) [58], which measures the average Euclidean distance between the predicted and ground truth flow vectors.
+The DSEC-Flow dataset (Gehrig et al., 2021a) consists of 24 challenging driving sequences with varying lighting conditions, fast motion, and occlusions.
+To assess our models on unseen data, we created a custom test set by splitting the training set, following the approach in Ponghiran et al.
+(2023).
+To evaluate the accuracy of the predicted optical flow, we use average end point error (AEE) (Zhu et al., 2018a), which measures the average Euclidean distance between the predicted and ground truth flow vectors.
 $\vspace{-1ex}AEE=\frac{1}{n}\sum_{n}{\lVert{(u,v)}_{pred}-{(u,v)}_{gt}\rVert}_{2}$ (5)
-For experiments on DSEC-flow [13], we use a fully spiking [29] and hybrid [36] multi-scale encoder-decoder network inspired by EV-FlowNet [58] as our backbone architecture.
+For experiments on DSEC-flow (Gehrig et al., 2021a), we use a fully spiking (Kosta & Roy, 2023) and hybrid (Negi et al., 2024) multi-scale encoder-decoder network inspired by EV-FlowNet (Zhu et al., 2018a) as our backbone architecture.
 We augment this architecture by integrating TSkips into the existing skip connections between the encoder and decoder blocks, as depicted in Fig.
@@
 Due to the inherent structure of our backbone, the search space for optimal TSkips configurations is relatively small, focusing primarily on selecting which existing skip connection to modify and determining the appropriate temporal delay ($\Delta t$).
-Figure 5: Fully spiking EV-FlowNet [58] architecture with forward and backward TSkips.
-Each TSkip replaces a pre-existing skip connection between the encoder and decoder.
+Figure 5: Fully spiking EV-FlowNet (Zhu et al., 2018a) architecture with forward and backward TSkips.
+Each TSkips replaces a pre-existing skip connection between the encoder and decoder.
 Although shown together for conciseness, forward and backward connections are evaluated independently.
@@
 DVS128-Gesture
-The IBM DVS128 Gesture dataset [21] contains 1342 instances of 11 hand gestures, captured with a DVS128 [33] camera.
-For experiments on DVS128 Gesture, we use a ResNet18 [17] backbone architecture [5] and explore various configurations by modifying its basic blocks.
+The IBM DVS128 Gesture dataset (Hu et al., 2022) contains 1342 instances of 11 hand gestures, captured with a DVS128 (Lichtsteiner et al., 2008) camera.
+For experiments on DVS128 Gesture, we use a ResNet18 (He et al., 2015) backbone architecture (Cai et al., 2020) and explore various configurations by modifying its basic blocks.
 This search includes modifying the input and output channels of each layer, kernel sizes, stride, and network depth.
-To integrate TSkips in the backbone, we remove the original ResNet18 skip connections and search for optimal TSkip placement and corresponding $\Delta t$.
-For optical flow estimation and gesture recognition, we introduce a learnable scaling factor, $\alpha$, within the TSkip to further enhance temporal processing (detailed in Section. 3.3).
+To integrate TSkips in the backbone, we remove the original ResNet18 skip connections and search for optimal TSkips placement and corresponding $\Delta t$.
+For optical flow estimation and gesture recognition, we introduce a learnable scaling factor, $\alpha$, within the TSkips to further enhance temporal processing (detailed in Section. 3.3).
 SHD and SSC
-The SHD [8] dataset comprises 10,000 recordings of spoken digits (zero to nine in English and German), while the larger SSC [8] dataset contains 100,000 recordings of spoken words from various speakers.
+The SHD (Cramer et al., 2022) dataset comprises 10,000 recordings of spoken digits (zero to nine in English and German), while the larger SSC (Cramer et al., 2022) dataset contains 100,000 recordings of spoken words from various speakers.
 These datasets present challenges for audio classification, requiring the network to capture and process subtle temporal patterns within spike trains that encode spoken sounds.
 We focus on the top-one classification task for all 35 classes in SSC and all 20 classes in SHD.
-For our experiments on speech recognition, we use a fully connected Multi Layer Perceptron(MLP) [47] as our backbone architecture.
+For our experiments on speech recognition, we use a fully connected Multi Layer Perceptron(MLP) (Rumelhart et al., 1986) as our backbone architecture.
 We explore various MLP configurations by adjusting the input and output features of each layer, the overall network depth, and the placement and $\Delta t$ of TSkips.
-In these experiments, we omit $\alpha$ used in the EV-FlowNet and DVS128 Gesture TSkip architectures.
-To improve generalization, we augment the training sets for DVS128 Gesture, SHD and SSC with channel jitter and random noise [49].
+In these experiments, we omit $\alpha$ used in the EV-FlowNet and DVS128 Gesture TSkips architectures.
+To improve generalization, we augment the training sets for DVS128 Gesture, SHD and SSC with channel jitter and random noise (Shen et al., 2023).
 We use sequence lengths (T) of 10, 30, and 99 for DSEC-flow, DVS128 Gesture, and SHD/SSC, respectively.
-All models are trained on an NVIDIA A40 GPU using the Adam optimizer [28].
-For DSEC-flow, we use a multi-step learning rate scheduler (initial rate: $10^{-3}$, scaled by 0.7 every 10 epochs) and the supervised mean squared error (MSE) loss [29, 36] for 200 epochs.
-For all other datasets, we use a Cosine Annealing Learning Rate Scheduler [34] (initial rate: 0.001, minimum rate: $5\times 10^{-6}$, updated every 10 iterations) for 100 epochs.
-Details on the constraints on the search parameters and TSkip configurations for DSEC-flow, DVS128 Gesture, SHD, and SSC can be found in the Supplementary material.
+All models are trained on an NVIDIA A40 GPU using the Adam optimizer (Kingma & Ba, 2014).
+For DSEC-flow, we use a multi-step learning rate scheduler (initial rate: $10^{-3}$, scaled by 0.7 every 10 epochs) and the supervised mean squared error (MSE) loss (Kosta & Roy, 2023; Negi et al., 2024) for 200 epochs.
+For all other datasets, we use a Cosine Annealing Learning Rate Scheduler (Loshchilov & Hutter, 2017) (initial rate: 0.001, minimum rate: $5\times 10^{-6}$, updated every 10 iterations) for 100 epochs.
+Details on the constraints on the search parameters and TSkips configurations for DSEC-flow, DVS128 Gesture, SHD, and SSC can be found in Appendix D.1.
 4.2 Ablation Study
-Figure 6: Ablation study results on the SHD dataset demonstrating the impact of varying the (a) temporal delay $(\Delta t)$, (b) TSkip position and (c) network depth on classification accuracy.
+Figure 6: Ablation study results on the SHD dataset demonstrating the impact of varying the (a) temporal delay $(\Delta t)$, (b) TSkips position and (c) network depth on classification accuracy.
 As detailed in Section 3.4, the search space of possible TSkips configurations is exponentially large.
-To effectively navigate this complex search space, we performed an ablation study on an 8-layer MLP baseline found using [27] on the SHD dataset.
-We varied the temporal delay ($\Delta t$), TSkip position and network depth using concatenation-based [22] TSkips for our analysis (see Supplementary for detailed baseline network configuration and comparisons with addition-based TSkips).
+To effectively navigate this complex search space, we performed an ablation study on an 8-layer MLP baseline found using Kim et al.
+(2022) on the SHD dataset.
+We varied the temporal delay ($\Delta t$), TSkips position and network depth using concatenation-based (Huang et al., 2017) TSkips for our analysis (see Appendix C.6 for detailed baseline network configuration).
 Impact of Temporal Delay $\mathbf{(\Delta t)}$:
@@
 This highlights the importance of optimizing $\Delta t$ relative to the sequence length ($T$), as both excessively small and large delays can hinder performance.
-Impact of TSkip Position:
-Next, we investigated the impact of TSkip position on classification accuracy (Fig.
+Impact of TSkips Position:
+Next, we investigated the impact of TSkips position on classification accuracy (Fig.
 6(b)).
@@
 This suggests that propagating temporal information to deeper layers is crucial for enhancing model performance.
-Note, the optimal placement varies based on forward or backward TSkip.
+Note, the optimal placement varies based on forward or backward TSkips.
 Impact of Network Depth:
@@
 This suggests the optimal configuration of TSkips and network depth are intertwined.
-These findings underscore the complex interplay between $\Delta t$, TSkip position (origin and destination layers), and network depth, which we set as our search parameters.
-For the results presented in the following sections, we leveraged [27] to identify optimal baseline and concatenation-based [22] TSkip architectures, constraining the architecture search by the number of parameters.
+These findings underscore the complex interplay between $\Delta t$, TSkips position (origin and destination layers), and network depth, which we set as our search parameters.
+For the results presented in the following sections, we leveraged Kim et al.
+(2022) to identify optimal baseline and concatenation-based (Huang et al., 2017) TSkips architectures (see Appendix C.5 for comparisons with addition-based TSkips), constraining the architecture search by the number of parameters.
 4.3 DSEC-Flow Results
-Incorporating TSkips into the fully spiking and hybrid EV-FlowNet architectures [29, 36], as detailed in Section 4.1, significantly lowers AEE on the DSEC-flow dataset (Table 1).
+Figure 7: Qualitative results on the DSEC-flow dataset.
+(a) Ground Truth (GT) mask, (b) Mini (3.4M) SNN baseline model mask, (c) Mini SNN (3.4M) with forward (F) TSkips, (d) Mini SNN (3.4M) with backward (B) TSkips, and (e) Base (13M) SNN baseline model mask.
+As observed, the Mini model that is 3.8$\times$ smaller than Base model performs just as well qualitatively when TSkips are incorporated.
+Incorporating TSkips into the fully spiking and hybrid EV-FlowNet architectures (Kosta & Roy, 2023; Negi et al., 2024), as detailed in Section 4.1, significantly lowers AEE on the DSEC-flow dataset (Table 1).
 Across both fully spiking and hybrid architectures, TSkips consistently yielded an average reduction in AEE of 14% and 9.5%, respectively.
 Notably, these improvements are achieved with models 3×3$\times$3 × smaller than baselines with comparable AEE.
-While SOTA methods like E-RAFT [14] and E-FlowFormer [32] achieve 6% and 10% lower AEE, respectively, these reductions come at the cost of increased complexity.
-E-RAFT, for example, while being 3×3$\times$3 × smaller than our best model, relies on complex mechanisms such as gated recurrent units and iterative flow refinement.
-Similalry, E-FlowFormer employs a computationally expensive transformer architecture and is trained on both a custom dataset and DSEC-flow.
-These complexities likely lead to higher inference energy compared to our lightweight TSkip architectures.
-Importantly, our approach achieves comparable qualitative performance to baselines 3×3$\times$3 × larger (Fig.
-7), highlighting its efficiency in enabling accurate temporal processing without increasing model complexity.
-Hyper parameters and a detailed analysis of inference energy for TSkips on DSEC-flow are in the Supplementary material.
-(a) GT
-(b) Mini - Baseline
-(c) Mini - F Tskip
-(d) Mini - B Tskip
-(e) Base - Baseline
-Figure 7: Qualitative results on the DSEC-flow dataset.
-(a) Ground Truth (GT) mask, (b) Mini (3.4M) SNN baseline model mask, (c) Mini SNN (3.4M) with forward (F) Tskip, (d) Mini SNN (3.4M) with backward (B) TSkip, and (e) Base (13M) SNN baseline model mask.
-As observed, the Mini model that is 3.8$\times$ smaller than Base model performs just as well qualitatively when TSkips are incorporated.
-Table 1: AEE on DSEC-flow (lower is better). Best performing SNN and hybrid baselines and forward(F)/backward(B) TSkip models are highlighted in bold.
+Although SOTA methods such as E-RAFT (Gehrig et al., 2021b) and E-FlowFormer (Li et al., 2023) achieve 6% and 10% lower AEE , respectively, these reductions come at the cost of increased complexity.
+E-RAFT, for example, while being 3×3$\times$3 × smaller than our best model, relies on complex processing of event data before its gated recurrent unit (GRU) update.
+These processing operations involve converting each event to a voxel grid, extracting features using CNNs and building a correlation volume.
+These dense kernel operations are computationally expensive during both training and inference, as they must be performed for every event sample.
+E-FlowFormer is a transformer-based architecture which relies on dense MAC operations in it’s self and cross attention blocks.
+These event processing steps lead to higher inference energy compared to TSkips architectures, which take advantage of the inherent sparsity and event-driven nature of SNNs, performing computationally cheaper sparse AC operations.
+Additionally, E-FlowFormer is trained on a large custom dataset and DSEC-flow, which introduces a significant increase in computation power and memory during training.
+Fig.
+7 provides a qualitative analysis of how the 3×3$\times$3 × larger baselines perform compared to the smaller TSkips variants which highlights the efficiency of TSkips in allowing accurate temporal processing without increasing model complexity.
+Hyper parameters and a detailed analysis of convergence and inference energy for TSkips on DSEC-flow models are in Appendix B.
+Table 1: AEE on DSEC-flow (lower is better). Best performing SNN and hybrid baselines and forward/backward TSkips models are highlighted in bold.
 Architecture #Params (M) AEE
 SNN Hybrid
-EV-FlowNet [14] 13.04 2.32 -
-LSTM-FlowNet [41] N/A [^1] N/A - not reported #Params 1.28 -
-E-RAFT [14] 5.3 0.79 -
-E-FlowFormer [32] N/A 0.76 -
+EV-FlowNet (Gehrig et al., 2021b) 13.04 2.32 -
+LSTM-FlowNet (Ponghiran et al., 2023) N/A [^1] N/A - not reported #Params 1.28 -
+E-RAFT (Gehrig et al., 2021b) 5.3 0.79 -
+E-FlowFormer (Li et al., 2023) N/A 0.76 -
 Our Models
 Base - Baseline 13.04 1.35 0.96
-Base + F TSkip 1.12 0.86
-Base + B TSkip 1.13 0.84
+Base + Forward TSkips 1.12 0.86
+Base + Backward TSkips 1.13 0.84
 Mini - Baseline 3.41 1.65 1.11
-Mini + F TSkip 1.44 1.02
-Mini + B TSkip 1.46 1.01
+Mini + Forward TSkips 1.44 1.02
+Mini + Backward TSkips 1.46 1.01
 Micro - Baseline 0.93 1.80 1.22
-Micro + F TSkip 1.57 1.12
-Micro + B TSkip 1.56 1.09
+Micro + Forward TSkips 1.57 1.12
+Micro + Backward TSkips 1.56 1.09
 Nano - Baseline 0.27 2.17 1.47
-Nano + F TSkip 1.86 1.37
-Nano + B TSkip 1.77 1.35
+Nano + Forward TSkips 1.86 1.37
+Nano + Backward TSkips 1.77 1.35
 Pico - Baseline 0.092 2.57 2.02
-Pico + F TSkip 2.19 1.78
-Pico + B TSkip 2.28 1.81
-Table 2: Classification accuracy on DVS-Gesture. Comparisons between forward(F)/backward(B) TSkips and SOTA models. Best models in bold.
+Pico + Forward TSkips 2.19 1.78
+Pico + Backward TSkips 2.28 1.81
+4.4 DVS128-Gesture, SHD and SSC Results
+Using TSkips in ResNet18 and MLP backbone architectures, as detailed in Section 4.1, consistently improved the classification accuracy of SNNs across the DVS Gesture, SHD, and SSC datasets (Tables 2 and 3).
+On the DVS Gesture dataset, incorporating TSkips resulted in a significant 8.37% improvement in accuracy over the baseline model, a convolutional SNN found by Kim et al.
+(2022).
+Importantly, our approach achieved this with significantly smaller models compared to SOTA, including a spiking transformer (Qin & Liu, 2024), demonstrating the efficiency of our proposed method.
+Furthermore, our method outperformed several other approaches, including spiking RNNs (Xing et al., 2020), methods incorporating learnable delays (Shrestha & Orchard, 2018), and methods using specialized LIF neurons or spatio-temporal feature extraction (Jiang & Zhang, 2024; Samadzadeh et al., 2023).
+This highlights the effectiveness of our approach in capturing and leveraging temporal dependencies for improved gesture recognition.
+Table 2: Classification accuracy on DVS-Gesture. Comparisons between forward/backward TSkips and SOTA models. Best models in bold.
 Method Base Model #Params (M) Accuracy (%)
-SLAYER [50] SNN (8 layer) N/A 93.6
-KLIF [23] VGG-11 N/A 93.75
-DECOLE [24] Custom N/A 95.2
-SNN-Skip [2] ResNet18 N/A 95.43
-SCRNN [54] Custom N/A 96.59
-STS-ResNet [48] ResNet18 N/A 96.7
-Mamba-Spike [42] Mambda 6.1 97.8
+SLAYER (Shrestha & Orchard, 2018) SNN (8 layer) N/A 93.6
+KLIF (Jiang & Zhang, 2024) VGG-11 N/A 93.75
+DECOLE (Kaiser et al., 2020) Custom N/A 95.2
+SNN-Skip (Benmeziane et al., 2023) ResNet18 N/A 95.43
+Streaming rollout SNN Kugele et al. (2021) DenseNet 0.8 95.56
+SCRNN (Xing et al., 2020) Custom N/A 96.59
+STS-ResNet (Samadzadeh et al., 2023) ResNet18 N/A 96.7
+Mamba-Spike (Qin & Liu, 2024) Mambda 6.1 97.8
 Our Models
 Baseline ResNet18 Backbone 0.30 88.75
-vRNN [^2] Backward skips are added to all baseline layers with $\Delta~{}t=1$. 0.30 89.12
-F TSkip 0.42 94.82
-B TSkip 0.38 95.97
-F + B TSkip 0.55 97.52
-4.4 DVS128-Gesture, SHD and SSC Results
-Using TSkips in ResNet18 and MLP backbone architectures, as detailed in Section 4.1, consistently improved the classification accuracy of SNNs across the DVS Gesture, SHD, and SSC datasets (Tables 2 and 3).
-On the DVS Gesture dataset, incorporating TSkips resulted in a significant 8.37% improvement in accuracy over the baseline model, a convolutional SNN found by [27].
-Importantly, our approach achieved this with significantly smaller models compared to SOTA, including a spiking transformer [42], demonstrating the efficiency of our proposed method.
-Furthermore, our method outperformed several other approaches, including spiking RNNs [54], methods incorporating learnable delays [50], and methods using specialized LIF neurons or spatio-temporal feature extraction [23, 48].
-This highlights the effectiveness of our approach in capturing and leveraging temporal dependencies for improved gesture recognition.
-On the SHD dataset, TSkips achieved an average accuracy gain of 8.6% over the 4-layer baseline (Baseline - 1) and 8.31% over the 8-layer baseline (Baseline - 2).
-Our approach also outperformed various recurrent architectures (vRNNs, spiking RNNs, and LSTMs) by substantial margins.
-Similarly, on the SSC dataset, TSkips yielded an average accuracy gain of 14% over Baseline - 1 and 16.04% over the 8-layer Baseline - 2, again surpassing the performance of recurrent architectures.
+vRNN [^2] Backward skips are added to all baseline layers with $\Delta t=1$. 0.30 89.12
+Forward TSkips 0.42 94.82
+Backward TSkips 0.38 95.97
+Forward + Backward TSkips 0.55 97.52
 Table 3: Classification accuracy on SHD and SSC datasets.
-Comparisons between forward(F), backward(B) and a combination (F+B) of both TSkips against SOTA models and two baseline models: a shallow 4-layer MLP network (Baseline - 1) and a deep 8-layer MLP (Baseline - 2).
+Comparisons between forward, backward and a combination of both TSkips against SOTA models and two baseline models: a shallow 4-layer MLP network (Baseline - 1) and a deep 8-layer MLP (Baseline - 2).
 Best models with a single and two TSkips in bold.
@@
 SOTA Models
-LSTM-LIF [56] 0.14 0.11 88.91 63.46
-SRNN [55] N/A N/A 90.4 74.2
-DL256-SNN-DLoss [51] 0.14 - 92.56 -
-SpikGRU [9] - 0.28 - 77
-radLIF [3] 3.9 3.9 94.62 77.40
-Spikingformer [57]333[57] was evaluated on SHD and SSC. [57] for learning rate, with minimal updates to the architecture. 1.98 1.99 82.68 72.43
-DCLS-Delays [52] 0.21 2.5 95.07 80.69
+LSTM-LIF (Zhang et al., 2023) 0.14 0.11 88.91 63.46
+SRNN (Yin et al., 2021) N/A N/A 90.4 74.2
+DL256-SNN-DLoss (Sun et al., 2023a) 0.14 - 92.56 -
+SpikGRU (Dampfhoffer et al., 2022) - 0.28 - 77.00
+radLIF (Bittar & Garner, 2022) 3.9 3.9 94.62 77.40
+Spikingformer (Zhou et al., 2023) [^3] Zhou et al. (2023) was evaluated on SHD and SSC with minimal updates to the architecture and tuned learning rate. 1.98 1.99 82.68 72.43
+DCLS-Delays (Sun et al., 2023b) 0.21 2.5 95.07 80.69
 Our Baseline MLPs
@@
 Our MLP Models with TSkips
-vRNN [^2] Backward skips are added to all baseline layers with $\Delta~{}t=1$. - 1 0.16 0.12 68.5 71.2
-F TSkip - 1 0.24 0.42 92.32 76.5
-B TSkip - 1 0.19 0.24 93.64 79.87
-F + B TSkip - 1 0.20 0.54 93.01 78.64
-vRNN [^2] Backward skips are added to all baseline layers with $\Delta~{}t=1$. - 2 1.04 1.04 71.35 72.24
-F TSkip - 2 1.12 1.08 94.15 78.98
-B TSkip - 2 1.16 1.14 93.86 79.65
-F + B TSkip - 2 1.28 1.42 94.73 80.23
-While TSkips achieves accuracy within 0.34% of SOTA model DCLS-Delays [16] on SHD and 0.46% on SSC, it offers several advantages: applicability across various network architectures, as shown in Section 4.3, efficiency due to very few new trainable parameters, and strong performance.
-Notably TSkips achieves comparable accuracy to [16] on the SSC dataset with a 44% smaller model.
+vRNN22footnotemark: 2 - 1 0.16 0.12 68.50 71.20
+Forward TSkips - 1 0.24 0.42 92.32 76.50
+Backward TSkips - 1 0.19 0.24 93.64 79.87
+Forward + Backward TSkips - 1 0.20 0.54 93.01 78.64
+vRNN22footnotemark: 2 - 2 1.04 1.04 71.35 72.24
+Forward TSkips - 2 1.12 1.08 94.15 78.98
+Backward TSkips - 2 1.16 1.14 93.86 79.65
+Forward + Backward TSkips - 2 1.28 1.42 94.73 80.23
+On the SHD dataset, TSkips achieved an average accuracy gain of 8.6% over the 4-layer baseline (Baseline - 1) and 8.31% over the 8-layer baseline (Baseline - 2).
+Our approach also outperformed various recurrent architectures (vRNNs, spiking RNNs, and LSTMs) by substantial margins.
+Similarly, on the SSC dataset, TSkips yielded an average accuracy gain of 14% over Baseline - 1 and 16.04% over the 8-layer Baseline - 2, again surpassing the performance of recurrent architectures.
+While TSkips achieves accuracy within 0.34% of SOTA model DCLS-Delays (Hammouamri et al., 2023) on SHD and 0.46% on SSC, it offers several distinct advantages.
+Specifically, the architectural constraint of DCLS-Delays to feedforward networks highlights a key area where TSkips provides a complementary alternative.
+Furthermore, TSkips as an architectural addition coupled with NAS allows for modifications to various network architectures, as shown in Section 4.3 and also offers efficiency due to very few new trainable parameters coupled with strong performance.
+In particular TSkips achieves accuracy comparable to Hammouamri et al.
+(2023) on the SSC dataset with a 44% smaller model.
 These improvements can be attributed to the ability of TSkips to capture long-term temporal dependencies, facilitated by the unrolled network structure, detailed in Section 3.
 Our method improves both the accuracy and convergence rate of the models by mitigating vanishing spikes and enabling more efficient gradient propagation during training.
-Analysis in the Supplementary material demonstrates faster convergence and lower energy consumption compared to standard SNNs and recurrent architectures.
-Additionally, we show that incorporatingTSkips in ANNs and hybrid models significantly improves gesture and speech classification.
+The analysis presented in Appendices A and C demonstrates lower energy consumption and faster convergence compared to standard SNNs and recurrent architectures.
+Additionally, we show that incorporatingTSkips in ANNs and hybrid models significantly improves speech classification (Appendix C.4).
+Figure 8: Trained network’s response to temporal patterns at inference, visualized by average spiking activity.
+For two randomly sampled SHD event streams ((a) and (b)), we compare the true event activity against that of a baseline MLP, a forward TSkips (FTSkips) network, the FTSkips network with the skip removed (F no skip), a backward TSkips (BTSkips) network, and the BTSkips network with the skip removed (B no skip).
+Sub-figures (c) and (d) show the average spiking activity specifically within the forward and backward TSkips, demonstrating spike generation aligned with the specified delay window.
+All plots are normalized by the 700-dimensional input feature representation.
+4.5 Observations
+To empirically validate our claim that TSkips enhance the network’s responsiveness to temporal patterns, we analyze the spike representations of trained models at inference.
+Fig.
+8 compares the average spiking activity of trained models with and without TSkips against the true event stream for two randomly sampled SHD data points.
+The TSkips models without a temporal skip have a zero tensor passing through the TSkips connection to ensure the model being evaluated does not undergo any modifications.
+Figures 8(a) and 8(b) show that the average spiking pattern of the trained models (even after removing the TSkips, called F/B no skip) aligns more closely with the ground truth average spike distribution compared to the baseline network.
+Removing TSkips during inference reduces the average spike rate and memory requirements, and the qualitative results validate that the TSkips model has learned the input event distribution and captures the input dynamics better than the baseline.
+We attribute this to TSkips aiding the adaptive LIF parameters and weights to learn more effectively during BPTT.
+Furthermore, Fig.
+8(c) and (d) isolate the average spiking activity within the forward (F) and backward (B) TSkips connections themselves.
+These plots reveal two key observations: first, spike generation within the TSkips aligns with the specified temporal delay $(\Delta t)$, and second, the activity mirrors a scaled version of the original input distribution.
+This provides strong empirical evidence that TSkips propagates relevant temporal information from the past.
+Furthermore, Figs.
+8(a) and (b) validate that TSkips improve learning rather than acting as additional connections.
+The distinct characteristics of F and B TSkips are also evident.
+FTSkips exhibit more distributed spike patterns and higher variability, likely due to processing denser latent representations from earlier network stages.
+The smaller and sparser activity in BTSkips might facilitate more nuanced error correction during BPTT.
+Crucially, both types of TSkips demonstrate sensitivity to the entire event stream, including the sparse tail end, as seen in Figures 8(a) and (b).
+The sustained activity in Fig.8(a) and (b) during the later, sparser parts of the sequence in both (TSkips/no TSkips) model evaluations at inference indicates integration of information from earlier, denser time steps.
+This ability of TSkips to integrate information across different temporal contexts allows the network to build richer latent representations.
+This mechanism is particularly important for mitigating vanishing gradients/spikes during BPTT, as TSkips establish direct pathways for information and gradients across long temporal distances, ensuring learning even when events are sparser over the sequence length.
+We provide quantitative results on these ‘no skip’ models for the DSEc-flow and SHD datasets in Appendixes B.2 and  C.2.
+4.6 Scalability of TSkips
+To assess the scalability of TSkips to deeper architectures and more complex visual tasks, we evaluated our method on the CIFAR10-DVS (Li et al., 2017) dataset.
+While the DSEC-flow results (Section 4.3) demonstrated the effectiveness of TSkips on a complex real-world task, the underlying EV-FlowNet architecture is custom-designed.
+To further validate that TSkips works on standard, widely adopted architectures in a more complex event-based setting, we applied our method to larger models like spiking ResNet18 and VGG11, incorporating the NDA data augmentation technique (Li et al., 2022b) as our baseline.
+The results presented in Table 4 consistently demonstrate that TSkips improve accuracy over baseline models in these deeper networks.
+Table 4: Classification accuracy on the CIFAR10-DVS dataset, comparing the performance of baseline spiking ResNet18 and VGG11 models against their TSkips-augmented counterparts, illustrating the scalability of our approach.
+We use the NDA (Li et al., 2022b) spiking ResNet18 and VGG11 models as our baselines.
+Method Base Model #Params (M) Accuracy (%)
+SOTA Models
+Spikingformer (Zhou et al., 2023) Single-stage ViT 2.57 81.3
+$\alpha$-SSA (Xiao et al., 2025) Multi-stage ViT 1.54 82.24
+SpikingResformer (Shi et al., 2024) ResNet+ViT 35.52 84.8
+Basline Models
+Baseline - 1 (NDA (Li et al., 2022b)) ResNet18 11.700 78.00
+Baseline - 2 (NDA (Li et al., 2022b)) VGG11 132.86 81.70
+TSkips Models
+FTSkips - 1 ResNet18 11.708 81.08
+BTSkips - 1 ResNet18 12.224 82.93
+FTSkips - 2 VGG11 134.04 82.56
+BTSkips - 2 VGG11 134.04 83.01
+The CIFAR10-DVS dataset (Li et al., 2017) provides a valuable test case, showing that TSkips can provide accuracy gains even for event data with artificially generated temporal information from repeated closed-loop movements.
+These results offer insight into the effectiveness of TSkips in challenging vision/audio tasks that possess rich temporal dynamics, as detailed in Sections 4.3 and 4.4, demonstrating their ability to improve performance in deeper large-scale networks on event data.
+However, it is important to consider that smaller transformer models, such as Zhou et al.
+(2023); Xiao et al.
+(2025); Shi et al.
+(2024) can achieve similar or better performance with significantly fewer parameters.
+Beyond specific task performance, the vast search space for optimal TSkips configurations (temporal delay $(\Delta t)$ and skip positions) is a recognized challenge.
+This motivated our use of training-free NAS for efficient exploration.
+As detailed in Section 3.4, the NASWOT-SAHD (Kim et al., 2022) proxy inherently guides this search by favoring networks with high linear separability in LIF activations across time steps, suggesting robustness beyond just favorable initializations.
+To confirm that TSkips consistently provide performance improvements over strong baselines identified by NAS, we analyzed the best-found baseline and their TSkips-augmented variants across three random seeds.
+These results for DSEC-flow (Gehrig et al., 2021a) are provided in Appendix D.2, Table 12, further validating the consistent improvement in inference with TSkips.
+5 Discussion
+While our approach demonstrates promising results, several avenues for future investigation remain.
+Enhancing the energy efficiency of discovered architectures is one key area, potentially through energy-based regularization within a training-free NAS proxy.
+Furthermore, mitigating the increase in memory usage associated with TSkips and temporal unrolling requires exploring memory-efficient training techniques or alternative architectures.
+Future research could also investigate making the temporal delays learnable/ dynamically changing and inference energy optimization techniques to further refine the efficiency of our method.
+Implementing TSkips in neuromorphic or biological systems presents distinct challenges.
+The storage of historical neural states to enable connections across time steps necessitates memory allocation, which can be a significant constraint on neuromorphic hardware with limited on-chip resources.
+TSkips can be implemented on standard hardware by storing and retrieving past layer states, incurring memory and energy costs for simulation.
+The analysis in Sec.
+4.5 provides a promising solution to this problem, where the trained models can be deployed without hidden state storage and passing a null input through the TSkips.
+On neuromorphic hardware, TSkips can be realized through direct routing with delays, where information from an earlier layer is directly routed to a later layer with a controlled temporal delay.
+This can be achieved using physical delay lines or digital buffers within the chip’s architecture, potentially offering energy efficiency by bypassing computations in intermediate layers.
+Importantly, our findings in Section 5 show that trained TSkips models can perform well even without explicitly using the temporal skip at inference, reducing hardware memory and energy requirements.
+Biologically, analogies might be found in short-term memory mechanisms or axonal delays that could bridge temporal gaps in information flow.
+Overcoming connectivity and communication latency limitations in neuromorphic hardware will also be crucial for realizing flexible temporal skips.
+Balancing the performance gains of TSkips with these hardware costs will guide future work exploring memory-efficient encoding of historical states, incorporating hardware constraints into the NAS process, and investigating how to mitigate the inherent hardware delays using TSkips.
 6 Conclusion
@@
 This work was supported by, the Center for the Co-Design of Cognitive Systems (COCOSYS), a DARPA-sponsored JUMP 2.0 center, the Semiconductor Research Corporation (SRC), and the National Science Foundation.
-Supplementary Material
-6 Inference Energy Estimation
-To validate the observation that TSkips exhibit reduced energy consumption compared to baseline SNNs and standard recurrent methods (vRNNs [47] and LSTMs [19]) for various tasks (Sections 4.3 and 4.4), we estimate inference energy.
-These estimations are made using the method from [46, 30], which considers the energy required for sparse accumulate (AC) operations in SNNs ($E_{AC}=0.9pJ$) and dense multiply-and-accumulate (MAC) operations in ANNs ($E_{MAC}=4.6pJ$) based on a 45nm CMOS technology [20].
+Appendix A Inference Energy Estimation
+To validate the observation that TSkips exhibit reduced energy consumption compared to baseline SNNs and standard recurrent methods (vRNNs (Rumelhart et al., 1986) and LSTMs (Hochreiter, 1997)) for various tasks (Sections 4.3 and 4.4), we estimate inference energy.
+These estimations are made using the method from Rueckauer et al.
+(2017); Lee et al.
+(2021), which considers the energy required for sparse accumulate (AC) operations in SNNs ($E_{AC}=0.9pJ$) and dense multiply-and-accumulate (MAC) operations in ANNs ($E_{MAC}=4.6pJ$) based on a 45nm CMOS technology (Horowitz, 2014).
 We calculated the total inference energy ($E_{total}$) as:
@@
 Here, $T$ is the sequence length, $\mathcal{N}$ is the number of neurons in the layer, $\mathcal{C}$ is the number of synaptic connections per neuron, and $\mathcal{M}$ is the average spike rate over $T$.
-6.1 Inference Energy Ablation Study
-Figure 8: Ablation study on the SHD dataset demonstrating the impact of varying the (a) temporal delay $(\Delta t)$, (b) TSkip position and (c) network depth on inference energy.
-Building upon our ablation study in Section 4.2, we further investigated the impact of different TSkip configurations on inference energy, estimated using Eq.
+A.1 Inference Energy Ablation Study
+Figure 9: Ablation study on the SHD dataset demonstrating the impact of varying the (a) temporal delay $(\Delta t)$, (b) TSkips position and (c) network depth on inference energy.
+Building upon our ablation study in Section 4.2, we further investigated the impact of different TSkips configurations on inference energy, estimated using Eq.
 6.
-Specifically, we explored the impact of varying temporal delay ($\Delta t$), TSkip position, and network depth on an 8-layer multilayer perceptron (MLP) baseline model with concatenation-based TSkips, found using the Neural Architecture Search (NAS) framework [27].
+Specifically, we explored the impact of varying temporal delay ($\Delta t$), TSkips position, and network depth on an 8-layer multilayer perceptron (MLP) baseline model with concatenation-based TSkips, found using the Neural Architecture Search (NAS) framework (Kim et al., 2022).
 This analysis is crucial because TSkips increases the average spike rate within the network, potentially leading to higher inference energy consumption.
@@
 We investigated the impact of varying $\Delta t$ for both forward and backward TSkips on inference energy (Fig.
-8(a)).
+9(a)).
 Forward connections originated from the input data and terminated at the second layer, while backward connections originated from the final output and terminated at the seventh layer.
@@
 Next, we investigated the impact of varying TSkips placement on inference energy (Fig.
-8(b)).
+9(b)).
 With a fixed $\Delta t=16$, we varied the destination layer of both forward (originating from the input data) and backward (originating from the final output) connections.
@@
 Finally, we analyzed the impact of TSkips on inference energy across models of varying depth (Fig.
-8(c)).
+9(c)).
 Forward connections originated from the input, and backward connections originated from the final output, with $\Delta t=16$.
 Our analysis revealed that increasing the number of layers led to a significant increase in inference energy.
-This analysis of inference energy with varying TSkip parameters further emphasizes the complex interplay between $\Delta t$, TSkip position, and network depth.
-It highlights the importance of using methods like [27] to find optimal network architectures and TSkip configurations.
-While [27] does not explicitly optimize for minimal energy consumption, it prioritizes networks with diverse spiking activity by quantifying the dissimilarity between spiking patterns.
+This analysis of inference energy with varying TSkips parameters further emphasizes the complex interplay between $\Delta t$, TSkips position, and network depth.
+It highlights the importance of using methods like Kim et al.
+(2022) to find optimal network architectures and TSkips configurations.
+While Kim et al.
+(2022) does not explicitly optimize for minimal energy consumption, it prioritizes networks with diverse spiking activity by quantifying the dissimilarity between spiking patterns.
 As a result, the identified configurations often utilize TSkips with larger $\Delta t$ values, enabling our networks to efficiently capture long-term patterns while maintaining low energy consumption.
-7 Additional Results - DSEC-flow
-Table 4: Comparison of inference energy $(E_{total})$ using Eq. 6 for fully spiking DSEC-flow models with and without forward (F) / backward (B) TSkips across all scales.
-Model #Params (×$\times 10^{6}$) $(\mathcal{N})$ Variant #OPSSNN (×$(\times 10^{9})$ Spike Rate $(\mathcal{M})$ $E_{total}$ (mJ)
+Appendix B Additional Results - DSEC-flow
+Table 5: Comparison of inference energy ($E_{total}$) using Eq. 6 for fully spiking DSEC-flow models with and without forward/backward TSkips across all scales.
+Model #Params (×$\times 10^{6}$) ($\mathcal{N}$) Variant #OPSSNN (×$\times 10^{9}$) Spike Rate ($\mathcal{M}$) $E_{total}$ (mJ)
 Base 13.04 Baseline 25.9 45.83 23.3
-F Tskip 32.7 57.8 29.4
-B Tskip 30.7 54.32 27.6
+Forward TSkips 32.7 57.8 29.4
+Backward TSkips 30.7 54.32 27.6
 Mini 3.41 Baseline 9.71 45.15 8.75
-F Tskip 11.7 54.23 10.5
-B Tskip 11.4 53.21 10.3
+Forward TSkips 11.7 54.23 10.5
+Backward TSkips 11.4 53.21 10.3
 Micro 0.93 Baseline 5.81 61.9 5.25
-F Tskip 6.14 65.4 5.52
-B Tskip 5.86 62.4 5.27
+Forward TSkips 6.14 65.4 5.52
+Backward TSkips 5.86 62.4 5.27
 Nano 0.27 Baseline 2.67 55.55 2.40
-F Tskip 2.81 58.47 2.53
-B Tskip 2.71 56.3 2.44
+Forward TSkips 2.81 58.47 2.53
+Backward TSkips 2.71 56.3 2.44
 Pico 0.092 Baseline 2.11 74.12 1.90
-F Tskip 2.29 80.6 2.07
-B Tskip 2.25 78.9 2.02
-7.1 Inference Energy Analysis
+Forward TSkips 2.29 80.6 2.07
+Backward TSkips 2.25 78.9 2.02
+B.1 Inference Energy Analysis
 As demonstrated in Section 4.3, TSkips reduce average endpoint error (AEE) without increasing model complexity or energy consumption for optical flow estimation on the DSEC-flow dataset.
@@
 6.
-Table 4 demonstrates the effectiveness of incorporating TSkips into fully spiking EV-FlowNet architectures [29].
+Table 5 demonstrates the effectiveness of incorporating TSkips into fully spiking EV-FlowNet architectures (Kosta & Roy, 2023).
 By efficiently utilizing temporal information, our approach achieves comparable AEE to baseline models 3×3$\times$3 × larger, while significantly reducing inference energy.
@@
 This highlights the potential of our method for accurate, robust, and efficient optical flow prediction.
-7.2 Convergence Analysis
-Figure 9: Convergence analysis of fully spiking [29] and hybrid [36] EV-FlowNet architectures on the DSEC-flow dataset.
-The plots compare the convergence behavior of baseline models (Base and Nano) with their smaller forward (F) and backward (B) TSkip-augmented counterparts (Mini and Pico, respectively).
-As discussed in Section 3, incorporating TSkips can lead to easier training and faster convergence for both fully spiking [29]and hybrid [36] EV-FlowNet architectures.
+B.2 Inference Accuracy Analysis
+As detailed in Sec.
+4.5 the trained TSkips models were evaluated at inference after removing the temporal hidden state propagation through the TSkips.
+We provide qualitative results for these models in Table 6 and see that the DSEC models retain performance even without explicitly storing temporal states at inference.
+We attribute this to the learnable scaling factor $(\alpha)$ within the TSkips as detailed in  3.3.
+This $(\alpha)$ dynamically determines the weighted contribution of both the current timestep’s information and the information from the delayed past timestep propagating through the skip connection.
+For our "no TSkips" inference analysis on DSEC, we used the learned $\alpha$ to control how much of the past hidden state’s contribution was zeroed out in the skip connection, allowing the model to still leverage the learned weighting of the current timestep’s information.
+Table 6: Comparison of AEE (lower is better) on DSEC-flow across baseline architectures, models with forward (F)/ backward TSkips and models no TSkips.
+Arch Baseline FTSkips F no TSkips BTSkips B no TSkips
+Base 1.35 1.12 1.13 1.13 1.13
+Micro 1.65 1.44 1.44 1.46 1.47
+Mini 1.80 1.57 1.58 1.56 1.56
+Nano 2.17 1.86 1.86 1.77 1.77
+Pico 2.57 2.19 2.20 2.28 2.30
+B.3 Convergence Analysis
+As discussed in Section 3, incorporating TSkips can lead to easier training and faster convergence for both fully spiking (Kosta & Roy, 2023)and hybrid (Negi et al., 2024) EV-FlowNet architectures.
 To validate this, we analyzed the convergence behavior of our DSEC-flow models across different scales.
-Figure 9 shows that both our largest (Base vs.
+Fig.
+10 shows that both our largest (Base vs.
 Mini with TSkips) and smallest (Nano vs.
@@
 This consistent trend across different model sizes highlights the benefits of TSkips in facilitating fast, efficient and effective training.
-7.3 Hyper parameters for DSEC-flow
-Table 5 specifies the TSkip configurations used to replace existing skip connections in the fully spiking [29] and hybrid [36] EV-FlowNet architectures.
-Here, “TSkip Position” indicates the targeted connection, with ‘1’ representing the longest skip (between the first encoder and last decoder blocks) and ‘3’ the shortest (between the last encoder and first decoder blocks), see Fig.
+Figure 10: Convergence analysis of fully spiking (Kosta & Roy, 2023) and hybrid (Negi et al., 2024) EV-FlowNet architectures on the DSEC-flow dataset.
+The plots compare the convergence behavior of baseline models (Base and Nano) with their smaller forward (F) and backward (B) TSkips -augmented counterparts (Mini and Pico, respectively).
+B.4 Hyper parameters for DSEC-flow
+Table 7 specifies the TSkips configurations used to replace existing skip connections in the fully spiking (Kosta & Roy, 2023) and hybrid (Negi et al., 2024) EV-FlowNet architectures.
+Here, “TSkips Position" indicates the targeted connection, with ‘1’ representing the longest skip (between the first encoder and last decoder blocks) and ‘3’ the shortest (between the last encoder and first decoder blocks), see Fig.
 5.
@@
 $\Delta t$ TSkip Pos $\Delta t$ TSkip Pos
-Base + F TSkip 3 1 3 1
-Base + B TSkip 4 1 5 1
-Mini + F TSkip 3 1 4 2
-Mini + B TSkip 3 2 3 1
-Micro + F TSkip 4 2 3 2
-Micro + B TSkip 3 3 3 3
-Nano + F TSkip 3 1 3 2
-Nano + B TSkip 4 2 3 2
-Pico + F TSkip 4 1 3 1
-Pico + B TSkip 3 2 4 2
-8 Additional Results - DVS128 Gesture, SHD and SSC
-8.1 Inference Energy Analysis
+Base + Forward TSkips 3 1 3 1
+Base + Backward TSkips 4 1 5 1
+Mini + Forward TSkips 3 1 4 2
+Mini + Backward TSkips 3 2 3 1
+Micro + Forward TSkips 4 2 3 2
+Micro + Backward TSkips 3 3 3 3
+Nano + Forward TSkips 3 1 3 2
+Nano + Backward TSkips 4 2 3 2
+Pico + Forward TSkips 4 1 3 1
+Pico + Backward TSkips 3 2 4 2
+Appendix C Additional Results - DVS128 Gesture, SHD and SSC
+C.1 Inference Energy Analysis
 As demonstrated in Section 4.4,TSkips improves classification accuracy for speech recognition compared to standard recurrent models (vRNNs and LSTMs).
 Additionally, we state that TSkips also reduce inference energy compared to these models.
-This energy efficiency is further confirmed by the estimations presented in Table 6 for both shallow (4-layer) and deep (8-layer) MLP [47] networks on the SHD [8] dataset.
+This energy efficiency is further confirmed by the estimations presented in Table 8 for both shallow (4-layer) and deep (8-layer) MLP (Rumelhart et al., 1986) networks on the SHD (Cramer et al., 2022) dataset.
 These estimations were calculated using Eq.
@@
 Table 8: Inference energy ($E_{total}$) best-performing models on the SHD dataset. Energy is calculated using Eq. 6.
-Model #Params (×$\times 10^{6}$) $(\mathcal{N})$OPS (×$(\times 10^{12})$ Spike Rate (×$\times 10^{4}$) $(\mathcal{M})$ $E_{\text{total}}$ (mJ)
+Model
+#Params
+(×$\times 10^{6}$)
+($\mathcal{N}$)
+#OPS
+(×$\times 10^{12}$)
+Spike Rate
+(×$\times 10^{4}$)
+($\mathcal{M}$)
+$E_{\text{total}}$
 Baseline - 1 0.16 0.0083 3.16 7.52
@@
 LSTM - 1 0.2 0.016 - 74.63
-F TSkip -1 0.24 0.038 5.62 34.79
-B TSkip - 1 0.19 0.022 4.95 20.37
-F + B TSkip - 1 0.20 0.032 7.42 29.06
+Forward TSkips -1 0.24 0.038 5.62 34.79
+Backward TSkips - 1 0.19 0.022 4.95 20.37
+Forward + Backward TSkips - 1 0.20 0.032 7.42 29.06
 Baseline - 2 1.04 0.815 7.63 734.26
@@
 LSTM - 2 1.1 1.196 - 5505.28
-F TSkip - 2 1.12 1.204 8.21 1083.90
-B TSkip - 2 1.16 1.213 8.23 1091.87
-F + B TSkip - 2 1.28 1.761 9.13 1585.70
-8.2 Convergence Analysis
+Forward TSkips - 2 1.12 1.204 8.21 1083.90
+Backward TSkips - 2 1.16 1.213 8.23 1091.87
+Forward + Backward TSkips - 2 1.28 1.761 9.13 1585.70
+C.2 Inference Accuracy Analysis
+Table 9: Comparison of accuracy on SHD for two baseline models, forward (F)/ backward TSkips and no TSkips.
+Model TSkips no TSkips
+Baseline - 1 84.32
+FTSkips - 1 92.32 90.71
+BTSkips - 1 93.64 92.78
+F+B TSkips - 1 93.01 91.76
+Baseline - 2 86.42
+FTSkips - 2 94.15 92.32
+BTSkips - 2 93.86 93.18
+F+B TSkips - 2 94.73 92.34
+As detailed in Sec.
+4.5 the trained TSkips models were evaluated at inference after removing the temporal hidden state propagation through the TSkips.
+We see that the SHD models without the TSkips have a slight decrease in inference accuracy.
+We attribute this slight decrease to less information being propagated to the later layers of the network through the TSkips.
+However, the models continue to perform better than the baselines, thus validating that TSkips helped to improve learning during training.
+C.3 Convergence Analysis
+Figure 11: Test accuracy convergence on the SHD dataset, comparing the best-performing models with TSkips (from Table 3) to a vRNN and an LSTM with  $0.2\times 10^{6}$ parameters.
 As discussed in Sections 3 and 4.4, incorporating TSkips can lead to easier training and faster convergence compared to recurrent models like vRNNs and LSTMs, as demonstrated in Fig.
@@
 These factors, combined with the ability of TSkips to capture long-term temporal dependencies, contribute to the improved training efficiency observed in our experiments.
-Figure 10: Test accuracy convergence on the SHD dataset, comparing the best-performing models with TSkips (from Table 3) to a vRNN and an LSTM with $0.2\times 10^{6}$ parameters.
-8.3 TSkips in ANNs and Hybrid Models
+C.4 TSkips in ANNs and Hybrid Models
 As discussed in Section 4.4, incorporating TSkips into ANNs and hybrid ANN-SNN models can lead to improved classification accuracy.
 Table 10 presents these results, demonstrating the effectiveness of TSkips in enhancing both ANN and hybrid model performance, surpassing the performance of vRNNs and LSTMs, respectively.
-Table 7: Classification accuracy of ANN and Hybrid ANN-SNN models on the SHD and SSC datasets. We compare the 4-layer MLP baseline (Baseline-1) with and without Forward (F) and Backward (B) TSkips.
+Table 10: Classification accuracy of ANN and Hybrid ANN-SNN models on the SHD and SSC datasets. We compare the 4-layer MLP baseline (Baseline-1) with forward/backward TSkips.
 Method #Params (×$(\times 10^{6})$ Accuracy (%)
@@
 vRNN 0.16 60.51 67.45
-F TSkip 0.24 69.08 84.54
-B TSkip 0.19 73.45 86.02
-F + B TSkip 0.20 73.25 85.74
+Forward TSkips 0.24 69.08 84.54
+Backward TSkips 0.19 73.45 86.02
+Forward + Backward TSkips 0.20 73.25 85.74
 LSTM 0.20 80.54 -
@@
 vRNN 0.12 68.29 71.43
-F TSkip 0.42 71.19 73.48
-B TSkip 0.24 70.23 72.87
-F + B TSkip 0.54 71.09 74.96
+Forward TSkips 0.42 71.19 73.48
+Backward TSkips 0.24 70.23 72.87
+Forward + Backward TSkips 0.54 71.09 74.96
 LSTM 0.20 72.3 -
@@
 This improved gradient propagation can lead to more stable and efficient training, further contributing to the enhanced performance observed in ANNs and hybrid models.
-8.4 Comparison between TSkip Operators
-To further analyze the impact of different TSkip connection types, we compared the performance of concatenation-based and addition-based TSkips on the 4-layer and 8-layer baselines (Table 3).
+C.5 Comparison between TSkip Operators
+To further analyze the impact of different TSkips connection types, we compared the performance of concatenation-based and addition-based TSkips on the 4-layer and 8-layer baselines (Table 3).
+Table 11: Comparison of classification accuracy on SHD and SSC datasets with concatenation-based (C) and addition-based (A) forward/backward TSkips on a 4-layer (Baseline -1) and 8-layer (Baseline - 2) MLP network.
+Method #Params (×$(\times 10^{6})$ Accuracy (%)
+SHD SSC SHD SSC
+Baseline - 1
+Baseline 0.16 0.12 84.32 64.19
+Forward TSkips - C 0.24 0.42 92.32 76.5
+Backward TSkips - C 0.19 0.24 93.64 79.87
+Forward + Backward TSkips - C 0.20 0.54 93.01 78.64
+Forward TSkips - A 0.19 0.37 87.65 72.13
+Backward TSkips - A 0.11 0.21 87.54 73.15
+Forward + Backward TSkips - A 0.11 0.40 88.13 75.12
+Baseline - 2
+Baseline 1.04 1.04 86.42 67.54
+Forward TSkips - C 1.12 1.08 94.15 78.98
+Backward TSkips - C 1.16 1.14 93.86 79.65
+Forward + Backward TSkips - C 1.28 1.42 94.73 80.23
+Forward TSkips - A 0.99 1.05 86.39 72.15
+Backward TSkips - A 0.97 1.01 87.32 75.53
+Forward + Backward TSkips - A 0.98 1.14 89.45 77.12
 Table 11 demonstrates the superior performance of concatenation-based TSkips over addition-based on the baseline models.
@@
 Furthermore, concatenation preserves information from both sources, preventing potential information loss that can occur with addition.
-Table 8: Comparison of classification accuracy on SHD and SSC datasets with concatenation-based (C) and addition-based (A) Forward (F) and Backward (B) TSkips on a 4-layer (Baseline -1) and 8-layer (Baseline - 2) MLP network.
-Method #Params (×$(\times 10^{6})$ Accuracy (%)
-SHD SSC SHD SSC
-Baseline - 1
-Baseline 0.16 0.12 84.32 64.19
-F TSkip - C 0.24 0.42 92.32 76.5
-B TSkip - C 0.19 0.24 93.64 79.87
-F + B TSkip - C 0.20 0.54 93.01 78.64
-F TSkip - A 0.19 0.37 87.65 72.13
-B TSkip - A 0.11 0.21 87.54 73.15
-F + B TSkip - A 0.11 0.40 88.13 75.12
-Baseline - 2
-Baseline 1.04 1.04 86.42 67.54
-F TSkip - C 1.12 1.08 94.15 78.98
-B TSkip - C 1.16 1.14 93.86 79.65
-F + B TSkip - C 1.28 1.42 94.73 80.23
-F TSkip - A 0.99 1.05 86.39 72.15
-B TSkip - A 0.97 1.01 87.32 75.53
-F + B TSkip - A 0.98 1.14 89.45 77.12
-8.5 Hyper parameters for DVS128 Gesture, SHD and SSC
-Table 9 presents the top-ranked network architectures and TSkip configurations identified by [27], with corresponding results shown in Tables  2 and 3.
+C.6 Hyper parameters for DVS128 Gesture, SHD and SSC
+Table 13 presents the top-ranked network architectures and TSkips configurations identified by NASWOT-SAHD (Kim et al., 2022), with corresponding results shown in Tables  2 and 3.
 The table specifies the following for each configuration:
@@
       SHD and SSC: Specifies the number of input channels for each MLP layer, including the final fully connected layer.
-Channels represented as $n*2$ indicate that concatenation-based TSkips were used, doubling the number of channels at that layer.
+The channels represented as $n*2$ indicate that concatenation-based TSkips were used, doubling the number of channels $(n)$ in that layer.
 This representation is used to highlight that the results shown in Table 11, which uses addition-based TSkips, are based on the same underlying network architectures.
   2. 2.
-    TSkip parameters: The temporal delay ($\Delta t$) and the origin and destination layers $(from\rightarrow to)$ for Forward (F), Backward (B), and combined (F+B) TSkips.
+    TSkips parameters: The temporal delay ($\Delta t$) and the origin and destination layers $(from\rightarrow to)$ for forward (F), backward (B) and combined (F + B) TSkips.
   3. 3.
@@
     1. (a)
-      Leak and threshold initialization values for the adaptive LIF neurons: 0.6 and 15, respectively.
+      Leak and threshold initialization values for adaptive LIF neurons: 0.6 and 15, respectively.
     2. (b)
       Dataset-specific dropout rates: 0.4 for SHD and 0.2 for SSC.
-Table 9: Optimal Parameters for Training MLP Networks with Temporal Skip Connections on SHD and SSC Datasets.
-Variant Network Delay ($\Delta t)$ TSkip Position (from $\rightarrow$ to)
-DVS128 Gesture
-Baseline 2$\times$64$\times$64-3c80s1-3c80s1- 5c86s1-5c64s1-1c64s1-1c32s11 - -
-F TSkip 2$\times$64$\times$64-3c83s1-3c83s1-5c110s1- 3c94s1-1c94s1-1c32s11 5 1 $\rightarrow$ 2
-B TSkip 2$\times$64$\times$64-3c64s1-5c64s1- 5c64s1-5c76s1-1c76s1-1c32s11 5 4 $\rightarrow$ 3
-F + B TSkip 2$\times$64$\times$64-3c122s1-5c122s1- 3c122s1-5c64s1-1c64s1+1c32s11 F: 6 B: 8 F: 1 $\rightarrow$ 3 B: 5 $\rightarrow$ 4
-SHD
-Baseline - 1 700-124-288-144-20 - -
-F TSkip - 1 700-124*2-432-115-20 24 1 $\rightarrow$ 2
-B TSkip - 1 700*2-124-115-96-20 14 4 $\rightarrow$ 1
-F + B TSkip - 1 700*2-124-144-72*2-20 F: 18 B: 14 F: 2 $\rightarrow$ 4 B: 2 $\rightarrow$ 1
-Baseline - 2 700-512-288-224- 192-896-288-20 - -
-F TSkip - 2 700-532-288-524- 192-448*2-288-20 17 2 $\rightarrow$ 6
-B TSkip - 2 700-321-558*2-334 292-452-328-20) 12 3 $\rightarrow$ 5
-F + B TSkip - 2 700-321-228*2-334- 492*2-462-448-20 F: 10 B: 13 F: 2 $\rightarrow$ 5 B: 7 $\rightarrow$ 3
-SSC
-Baseline - 1 700-124-148-130-35 - -
-F TSkip - 1 700-324-298*2-145-35 15 1 $\rightarrow$ 3
-B TSkip - 1 700-184*2-179-230-35 12 3 $\rightarrow$ 2
-F + B TSkip - 1 700-364*2-257*2-195-35 F: 14 B: 8 F: 1 $\rightarrow$ 3 B: 4 $\rightarrow$ 2
-Baseline - 2 700-731-410-340- 150-184-58-35 - -
-F TSkip - 2 700-646-410-400- 250-184*2-135-35 17 2 $\rightarrow$ 6
-B TSkip - 2 700-663-373*2-337- 292-182-135-35 15 7 $\rightarrow$ 3
-F + B TSkip - 2 700-674-493*2-347*2- 278-192-89-35 F: 9 B: 11 F: 1 $\rightarrow$ 2 B: 7 $\rightarrow$ 3
-9 NASWOT-SAHD Correlation on TSkips
-To validate the effectiveness of NASWOT-SAHD [27] in identifying optimal network configurations, we analyzed the correlation between the accuracy and score of the identified networks, as shown in Fig.
-11.
-The scores are based on the Sparsity Aware Hamming Distance (SAHD) [27], and the correlation is measured using Kendall’s $\tau$.
-Our baseline models achieved a $\tau$ correlation of 0.63, consistent with the results reported in [27].
-However, the forward and backward TSkip network searches resulted in slightly lower correlations of 0.58 and 0.55, respectively.
-All results reported are for the top-ranked networks identified by  [27].
+Appendix D NASWOT-SAHD Validation
+D.1 Correlation on TSkips
+Figure 12: Correlation between accuracy and SAHD score for different network configurations (baseline, forward TSkips, and backward TSkips).
+To validate the effectiveness of NASWOT-SAHD (Kim et al., 2022) in identifying optimal network configurations, we analyzed the correlation between the accuracy and score of the identified networks, as shown in Fig.
+12.
+The scores are based on the Sparsity Aware Hamming Distance (SAHD) (Kim et al., 2022), and the correlation is measured using Kendall’s $\tau$.
+Our baseline models achieved a $\tau$ correlation of 0.63, consistent with the results reported in Kim et al.
+(2022).
+However, the forward and backward TSkips network searches resulted in slightly lower correlations of 0.58 and 0.55, respectively.
+All results reported are for the top-ranked networks identified by NASWOT-SAHD  (Kim et al., 2022).
 To ensure a focused and efficient search, we introduced several constraints:
@@
   2. 2.
-    TSkip Position: We constrained connections to not be within the same layer, ensuring that all connections originated from a different layer than the destination.
+    TSkips Position: We constrained connections to not be within the same layer, ensuring that all connections originated from a different layer than the destination.
   3. 3.
@@
     1. (a)
-      DSEC-Flow: No constraint was applied as we directly replaced existing skip connections in the fully spiking[29] and hybrid [36] EV-FlowNet architecture with TSkips, maintaining the original model size.
+      DSEC-Flow: No constraint was applied as we directly replaced existing skip connections in the fully spiking(Kosta & Roy, 2023) and hybrid (Negi et al., 2024) EV-FlowNet architecture with TSkips, maintaining the original model size.
     2. (b)
@@
       SHD and SSC (Baseline-2): $1.3\times 10^{6}$ parameters
-Figure 11: Correlation between accuracy and SAHD score for different network configurations (baseline, forward TSkip, and backward TSkip).
+D.2 Robustness across Random Seeds
+To validate the robustness of our NAS-identified configurations and the consistent performance enhancement provided by TSkips, we conducted additional experiments on the DSEC-flow (Gehrig et al., 2021a) dataset.
+Table 12 presents the Average Endpoint Error (AEE) for both baseline models and their TSkips-augmented variants, averaged across three different random seeds.
+These results, reported as mean $(\mu)\pm$ standard deviation $(\sigma)$, demonstrate the stability and reliable improvement offered by our approach.
+Table 12: DSEC flow results on performing NAS search for TSkips with multiple random seeds.
+Presented below is the AEE (mean$(\mu)$$\pm$ standard deviation$(\sigma)$) (lower is better) of the models found across 3 random seeds.
+Model $\Delta t$ Position AEE ($\mu\pm\sigma$)
+Base - Baseline - - 1.38 $\pm$ 0.09
+Base - FTSkip 3,2,3 1,1,2 1.16 $\pm$ 0.10
+Base - BTSkip 4,3,4 1,2,1 1.14 $\pm$ 0.13
+Mini - Baseline - - 1.72 $\pm$ 0.07
+Mini - FTSkip 3,2,2 1,1,2 1.49 $\pm$ 0.06
+Mini - BTSkip 3,2,3 2,1,2 1.47 $\pm$ 0.07
+Micro - Baseline - - 1.97 $\pm$ 0.11
+Micro - FTSkip 4,3,3 2,1,2 1.66 $\pm$ 0.10
+Micro - BTSkip 3,2,2 3,2,3 1.69 $\pm$ 0.15
+Nano - Baseline - - 2.23 $\pm$ 0.17
+Nano - FTSkip 3,3,2 1,1,2 1.85 $\pm$ 0.07
+Nano - BTSkip 4,3,3 2,1,2 1.78 $\pm$ 0.13
+Pico - Baseline - - 2.64 $\pm$ 0.08
+Pico - FTSkip 4,2,3 1,2,1 2.32 $\pm$ 0.13
+Pico - BTSkip 3,2,3 2,1,1 2.28 $\pm$ 0.17
+Table 13: Optimal Parameters for Training MLP Networks with Temporal Skip Connections on SHD and SSC Datasets.
+Variant Network
+Delay
+($\Delta t$)
+TSkip
+Position
+(from $\rightarrow$ to)
+DVS128 Gesture
+Baseline
+2$\times$64$\times$64-3c80s1-3c80s1-
+5c86s1-5c64s1-1c64s1-1c32s11
+F TSkips
+2$\times$64$\times$64-3c83s1-3c83s1-5c110s1-
+3c94s1-1c94s1-1c32s11
+5 1 $\rightarrow$ 2
+B TSkips
+2$\times$64$\times$64-3c64s1-5c64s1-
+5c64s1-5c76s1-1c76s1-1c32s11
+5 4 $\rightarrow$ 3
+F + B
+TSkips
+2$\times$64$\times$64-3c122s1-5c122s1-
+3c122s1-5c64s1-1c64s1+1c32s11
+F: 6
+B: 8
+F: 1 $\rightarrow$ 3
+B: 5 $\rightarrow$ 4
+SHD
+Baseline - 1 700-124-288-144-20 - -
+F TSkips - 1 700-124*2-432-115-20 24 1 $\rightarrow$ 2
+B TSkips - 1 700*2-124-115-96-20 14 4 $\rightarrow$ 1
+F + B
+TSkips
+- 1
+700*2-124-144-72*2-20
+F: 18
+B: 14
+F: 2 $\rightarrow$ 4
+B: 2 $\rightarrow$ 1
+Baseline - 2
+700-512-288-224-
+192-896-288-20
+F TSkips - 2
+700-532-288-524-
+192-448*2-288-20
+17 2 $\rightarrow$ 6
+B TSkips - 2
+700-321-558*2-334
+292-452-328-20)
+12 3 $\rightarrow$ 5
+F + B
+TSkips
+- 2
+700-321-228*2-334-
+492*2-462-448-20
+F: 10
+B: 13
+F: 2 $\rightarrow$ 5
+B: 7 $\rightarrow$ 3
+SSC
+Baseline - 1 700-124-148-130-35 - -
+F TSkips - 1 700-324-298*2-145-35 15 1 $\rightarrow$ 3
+B TSkips - 1 700-184*2-179-230-35 12 3 $\rightarrow$ 2
+F + B
+TSkips
+- 1
+700-364*2-257*2-195-35
+F: 14
+B: 8
+F: 1 $\rightarrow$ 3
+B: 4 $\rightarrow$ 2
+Baseline - 2
+700-731-410-340-
+150-184-58-35
+F TSkips - 2
+700-646-410-400-
+250-184*2-135-35
+17 2 $\rightarrow$ 6
+B TSkips - 2
+700-663-373*2-337-
+292-182-135-35
+15 7 $\rightarrow$ 3
+F + B
+TSkips
+- 2
+700-674-493*2-347*2-
+278-192-89-35
+F: 9
+B: 11
+F: 1 $\rightarrow$ 2
+B: 7 $\rightarrow$ 3